### v1

# Reasoning Efficiently Through Adaptive Chain-of-Thought Compression: A Self-Optimizing Framework

Abstract.
The introduction of Chain-of-Thought (CoT) reasoning has significantly enhanced the problem-solving capabilities of Large Language Models (LLMs) by prompting them to articulate intermediate reasoning steps rather than producing answers directly. This structured approach has led to significant improvements in accuracy and robustness across benchmarks in arithmetic, logic, and commonsense reasoning. Although effective, it introduces computational overhead. Each reasoning step requires additional token generation, which raises inference latency, memory usage, and key-value cache demands. These issues are particularly critical in domains such as software engineering, where deterministic logic and concise outputs are preferred.
To understand how CoT length impacts both model performance and user experience, we conduct an empirical study on widely used code generation datasets. We find that longer CoT does not always improve performance. Across benchmarks, excessively long reasoning often results in truncation, reduced accuracy, and latency increases of up to five times, with failed outputs consistently longer than successful ones. These findings challenge the assumption that longer reasoning is inherently beneficial and underscore the need for adaptive CoT control. Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive and self-enhancing framework for CoT compression that improves the conciseness of reasoning while preserving output quality, without relying on external compression tools. SEER follows an adaptive self-enhancing paradigm, where CoT responses are refined by combining Best-of-N sampling with task-aware adaptive filtering. The adaptive filter dynamically adjusts thresholds based on the token length of pre-inference model outputs, effectively controlling verbosity while reducing computational costs and preserving accuracy. We evaluate SEER on three software engineering tasks and one mathematical reasoning task. On average, SEER reduces CoT length by 42.1% without compromising performance, and accuracy even improves due to reduced truncation. Moreover, it substantially lowers truncation rates and eliminates most infinite reasoning loops, thereby enhancing both efficiency and robustness. These results highlight SEER as a practical approach to making the use of CoT more efficient, enabling CoT-enhanced LLMs to operate effectively even in resource-constrained settings and on complex tasks.
Chain-of-Thought (CoT), Large Language Models (LLMs), CoT Compression, Software Engineering, Reasoning Efficiency
copyright: none
1. Introduction
The introduction of Chain-of-Thought (CoT) marks a pivotal advancement in enhancing the reasoning capabilities of Large Language Models (LLMs), particularly for complex, multi-step tasks. By encouraging models to explicitly articulate intermediate reasoning steps, CoT enables a systematic, step-by-step breakdown of problems rather than merely producing an answer in a single attempt (Kojima et al., 2022; Wei et al., 2022). While CoT has demonstrated notable improvements in accuracy and robustness across tasks involving arithmetic, logic, and commonsense reasoning (Nye et al., 2021; Wang et al., 2023; Zhou et al., 2022), its effectiveness becomes even more pronounced in software engineering tasks. Such tasks typically demand precise logical reasoning, structured multi-step problem solving, and transparent decision-making, all of which are well supported by CoT. Recent work shows that CoT can substantially improve performance in multiple software engineering tasks (e.g., code generation, program synthesis, and automated debugging) by improving the interpretability of the model’s reasoning and aligning it with human problem-solving processes (Yao et al., 2023; Li et al., 2025).
However, this benefit comes with substantial costs. As most contemporary LLMs are based on the Transformer architecture (Vaswani et al., 2017), the attention mechanism incurs quadratic computational complexity with respect to context length, while the KV cache grows linearly with context size. Consequently, the verbose decoding of CoT significantly increases inference latency and memory usage, with KV-cache requirements emerging as a key bottleneck (Agrawal et al., 2024; Ho et al., 2024). This inefficiency is particularly critical in software engineering tasks, where timely responses are essential for developer productivity and effective decision-making. Moreover, standard CoT often “overthinks” even simple code generation tasks, generating excessive reasoning tokens despite the existence of straightforward solutions. In more complex scenarios (e.g., repairing software vulnerabilities), this overthinking can impair performance as it increases the inference costs and latencies (Cuadron et al., 2025). Consequently, reducing the length of CoT reasoning without compromising model performance remains a key challenge.
Recent research indicates that compressing the CoT in training data and subsequently fine-tuning the model can effectively manage the length of the generated output while still maintaining accuracy (Kang et al., 2025; Munkhbat et al., 2025a; Xia et al., 2025). Methods such as TokenSkip allow LLMs to eliminate semantically redundant tokens, achieving significant reductions (e.g., from 313 to 181 tokens on the GSM-8K dataset) with minimal performance loss (Xia et al., 2025). Advanced frameworks like Adaptive GoGI-Skip, R1-Compress, and CoLaR have further improved these gains by reducing CoT lengths by 45%–80% while preserving accuracy, as well as delivering 1.6 to 2 times speedups in inference (Zhuang et al., 2025; Wang et al., 2025b; Tan et al., 2025).
However, most of these methods are evaluated on mathematical reasoning datasets and commonsense QA, without assessing their effectiveness on software engineering problems; consequently, they may achieve good performance on mathematical reasoning tasks but exhibit substantially weaker results in software engineering contexts. Moreover, most of these approaches use external compression tools to compress CoT, such as GPT-4 in C3oT (Kang et al., 2025) and LLMLingua in TokenSkip (Xia et al., 2025), which may lead to information loss. Such information loss may lead to “thought leap” and ultimately influence the performance of the model (Xu et al., 2025b). For software engineering tasks, these compression tools may change the logic of the code or even make the code unable to execute.
In this work, we conduct an empirical study to investigate the impact of excessively long CoT outputs on coding tasks using four open-source reasoning models. Our empirical study shows that excessively long CoT reasoning can harm the performance of LLMs. On HumanEval, small LLMs suffer substantial accuracy drops from truncation (e.g., 60.36% to 46.34% on the 1.5B model), while larger models show limited gains since failed outputs are often linked to longer reasoning. In addition, lengthy CoT raises latency by up to five times, underscoring the need to control reasoning length. Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive and self-enhancing framework for CoT compression that improves the conciseness of reasoning while preserving output quality, without relying on external compression tools.
SEER adopts a self-enhancing paradigm, in which the model generates its own CoT-augmented responses and refines them via a two-stage data refinement process: 1) Best-of-N (BoN) sampling, which selects the shortest correct reasoning trace among multiple candidates to ensure both correctness and conciseness, and 2) an adaptive CoT filtering mechanism that dynamically calibrates maximum reasoning length based on the statistical properties of each dataset. This design enables task-specific control over verbosity, mitigating over-compression risks while eliminating pathological repetition patterns. Notably, SEER fine-tuned for one domain (e.g., code generation) also demonstrates strong transferability, preserving concise reasoning in other domains without retraining.
We evaluate SEER on three diverse software engineering tasks and one mathematical reasoning benchmark under an 8K token context constraint. Experimental results show that SEER achieves an average CoT length reduction of 42.1% while preserving or improving accuracy across all tasks. Compared to baselines such as TokenSkip and Naive BoN, our method remains robust in complex software engineering tasks, where other approaches often fail due to instability or logical corruption. Moreover, SEER mitigates infinite reasoning loops, which occur when a model repeatedly generates the same fragment until the context limit is reached, by up to 97.7%, substantially reducing truncation rates. In summary, our main contributions are as follows:
  • We conduct an empirical study on HumanEval to examine how excessively long CoTs affect coding tasks. The results show that overlong reasoning often causes truncation, loops, and latency, ultimately reducing accuracy and stability, motivating the need for effective CoT length control.
  • To address this, we propose SEER, a lightweight CoT compression framework that learns concise reasoning patterns from self-generated outputs, effectively reducing truncation and infinite loops without external modules.
  • Extensive experiments across software engineering and mathematical reasoning tasks demonstrate that SEER achieves an average 42.1% reduction in CoT length while maintaining or even improving pass@1 accuracy.
  • Ablation and cross-domain studies further confirm that adaptive CoT filtering is the key driver of efficiency and accuracy, with BoN sampling providing complementary benefits, and that SEER generalizes well across different task domains.
Paper Organization. The remainder of this paper is structured as follows: Section 2 reviews CoT efficiency challenges, the overthinking phenomenon, and self-generated fine-tuning. Section 3 presents a motivating empirical study on the impact of CoT length. Section 4 details the SEER framework, including BoN sampling and adaptive filtering. Section 5 outlines the datasets, baselines, and evaluation metrics. Section 6 reports experimental findings and ablation results, addressing performance, generalizability, and loop mitigation. Section 7 discusses broader implications and limitations, and Section 8 reviews related literature. Finally, Section 9 concludes the paper and outlines future research directions.
2. Background
2.1. Chain-of-Thought Efficiency
LLMs have achieved remarkable success in complex reasoning tasks, largely due to advanced prompting techniques like CoT (Wei et al., 2022). CoT elicits multi-step reasoning by instructing the model to generate a series of intermediate steps before producing a final answer. This explicit articulation of the thought process allows the model to deconstruct complex problems, track its progress, and correct intermediate errors, leading to significant performance gains in tasks that demand logical deduction, arithmetic calculation, and symbolic reasoning.
However, the verbosity of CoT introduces significant computational overhead. Each reasoning step requires additional token generation, directly increasing inference latency and resource consumption, particularly for the KV-cache (Agrawal et al., 2024; Ho et al., 2024). This overhead poses a critical bottleneck for real-time applications or systems with constrained computational budgets. Consequently, a key research direction in CoT efficiency is to strike an optimal balance: retaining the reasoning benefits of CoT while minimizing its associated costs. Approaches in this area typically explore methods such as pruning redundant steps, learning to generate more concise reasoning paths (Kang et al., 2025; Xia et al., 2025), or selectively applying CoT based on task complexity (Wang et al., 2025a).
2.2. Overthinking Phenomenon in CoT
Beyond simple inefficiency, the relationship between CoT length and accuracy is not a straightforward “longer is better.” Prior work shows that longer reasoning does not always translate into higher accuracy (Xie et al., 2025). Instead, many studies identify an “overthinking” phenomenon. Accuracy often improves with additional reasoning at the early stages, since the model has more opportunity to explore possible solution paths. However, once the reasoning exceeds an optimal range, performance begins to decline as the model produces redundant or even erroneous content (Chen et al., 2024b; Jin et al., 2024; Cuadron et al., 2025). This decline is a direct result of overthinking, where excessive reasoning introduces noise, deviates from the correct reasoning path, and ultimately undermines accuracy (Chen et al., 2024a). Taken together, these findings highlight that the key challenge is to carefully control CoT length so as to reduce overthinking while preserving accuracy.
2.3. Self-Generated Data for Self-Supervised Fine-Tuning
Recent research has demonstrated that LLMs can bootstrap their own capabilities by generating candidate data, applying automated quality filtering, and then fine-tuning on the filtered subset. STaR illustrates this approach. The model generates CoT rationales and answers, discards incorrect outputs, and fine-tunes iteratively on the correct rationale–answer pairs, reaching reasoning performance comparable to models that are thirty times larger (Zelikman et al., 2022). Additionally, extensions such as Quiet-STaR generalize STaR by generating token-level rationales during continued pretraining, improving zero-shot reasoning on benchmarks like GSM8K and CommonsenseQA without explicit task fine-tuning (Zelikman et al., 2024). These methods collectively form a cohesive pipeline: self-generation, automatic filtering, and self-fine-tuning. This structure allows LLMs to independently improve their reasoning, domain knowledge, and generalization capabilities without the need for human-labeled data.
3. Empirical Study
Before introducing our proposed framework, we conduct an empirical study to investigate how the length of CoT reasoning influences the performance, efficiency, and stability of LLMs in software engineering scenarios. Our analysis primarily targets coding tasks, a core application area of software engineering where precise and executable reasoning is essential. This study provides a quantitative assessment of the challenges introduced by excessively long CoT outputs, thereby motivating the design of our method.
3.1. Settings
3.1.1. Datasets
Our empirical study is conducted on the HumanEval (et al., 2021), which is a widely used benchmark consisting of manually curated Python programming problems. Each problem contains a function signature, a docstring, and a set of test cases for evaluating functional correctness. Additionally, in the final part of our empirical study, we use a more complex dataset, Codeforces (Penedo et al., 2025). Codeforces comprises challenging algorithmic optimization problems collected from one of the most popular websites among competitive programmers.
3.1.2. LLMs
We employ the DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025a) series of models to conduct our empirical study. This choice is motivated by the fact that it is one of the most popular reasoning models in recent times. Moreover, it is open-source, which facilitates deployment. We conduct experiments on four LLMs with sizes of 1.5B, 7B, 14B, and 32B.
3.1.3. Evaluation Metrics
To assess the performance of CoT, we employ two metrics.
[…]
Accuracy. We use Pass@1 as the primary metric to evaluate correctness. This metric calculates the proportion of problems for which the model produces a correct solution on the first attempt, directly reflecting its practical effectiveness in solving tasks.
[…]
Average Token Length. We report the average number of tokens generated in model outputs. This metric reflects the verbosity of the reasoning process and helps assess whether excessively long outputs introduce redundancy or cause truncation. It also facilitates a direct comparison of output lengths between models with and without CoT reasoning.
3.1.4. Implementation Details
[…]
3.2. Empirical Findings
[…]
Truncation Leads to Performance Degradation. Table 1 demonstrates that excessively long CoT outputs introduce severe truncation issues, which directly limit performance improvements. For the smallest model (DeepSeek-R1-Distill-Qwen-1.5B), accuracy drops sharply, with the pass@1 score falling from 60.36% to 46.34%. This loss coincides with an extreme average CoT length of 7,645 tokens. The underlying cause is frequent truncation, which removes critical parts of the solution. In fact, 63 out of 164 problems are truncated on this model. As the model size increases, the truncation problem becomes less frequent, and moderate gains emerge. The 7B and 14B models show performance gains with CoT, but these improvements are limited by long outputs and truncation. The 32B model still experiences truncation in about 8.5% of cases, which keeps its pass@1 unchanged at 90.24% with or without CoT.
Table 1. HumanEval Results of Reasoning Models (With vs. Without CoT)
Models HumanEval (With CoT) HumanEval (Without CoT)
Pass@1 (%) Avg Tokens Trunc Loop Pass@1 (%) Avg Tokens Trunc Loop
DeepSeek-R1-Distill-Qwen-1.5B 46.34 7,645 63 62 60.36 513 0 0
DeepSeek-R1-Distill-Qwen-7B 80.49 4,210 24 22 70.12 474 0 0
DeepSeek-R1-Distill-Qwen-14B 89.63 3,687 16 14 84.76 444 0 0
DeepSeek-R1-Distill-Qwen-32B 90.24 3,340 14 12 90.24 474 0 0
[…]
Finding 1:  Excessive CoT length induces harmful truncation, especially in small-scale LLMs. Small-scale LLMs are more easily affected by long CoTs because their context is quickly filled. For larger models, truncation is less frequent but still reduces the benefit, showing that longer reasoning does not always improve results once the context limit is reached.
[…]
Reasoning loop results in truncation. We manually examine truncation cases to better understand their impact on model performance. Our analysis shows that the model often falls into repetitive reasoning loops, repeatedly generating the same fragment. For example, it may continually output phrases such as “Wait, if …” until the maximum context length is reached. This behavior consumes a large number of tokens and severely reduces efficiency. As shown in Table 1, reasoning loops account for most truncation cases, with an average rate of 94.0%. When CoT generation is disabled, this problem disappears entirely. Larger models alleviate the issue to some extent, but reasoning loops are not fully eliminated.
These observations suggest that while CoT can enhance reasoning, uncontrolled generation often leads to loops and truncation, which harm both efficiency and reliability. Therefore, controlling CoT length is crucial to preserve its benefits without incurring instability.
[…]
Finding 2:  Reasoning loops are the dominant reason for truncation. They waste tokens, reduce efficiency, and persist across model scales, disappearing only when CoT is disabled. This highlights looping as a critical issue in CoT reasoning.
[…]
Is longer always better? While prior work (Jin et al., 2024) suggests that increasing the length of CoT reasoning improves accuracy, our findings reveal a different pattern. We conduct experiments on one of the most challenging problems in HumanEval (findmyway, 2024), namely HumanEval/129. Since the 1.5B model achieves almost no correct solutions in our experiments (pass@1 is nearly zero), we exclude it from further analysis. For each of the other models, we generate 10,000 responses and examine the distribution of CoT lengths, as shown in Figure 1, with the maximum token length set to 16,384. Contrary to the common assumption that longer reasoning consistently leads to better performance (Jin et al., 2024), our analysis reveals a clear pattern across all three models. Failed outputs are frequently associated with longer CoT lengths, while successful outputs are more concentrated around shorter lengths. For instance, in the 7B model, the median CoT length of failed generations (9,489 tokens) surpasses that of successful generations (8,296 tokens), a pattern that persists in both the 14B and 32B models. This suggests that excessively long reasoning chains may introduce redundancy, amplify error propagation, or even trigger context truncation, ultimately degrading performance. These results highlight that effective reasoning does not simply require “more thinking”, but rather appropriate reasoning lengths that balance sufficiency and conciseness.
Figure 1. Distribution of CoT lengths for passed and failed cases across DeepSeek-R1-Distill-Qwen models (7B, 14B, and 32B) on HumanEval/129. The results show that successful outputs do not necessarily require longer CoT reasoning. An excessively long reasoning chain can instead lead to a decline in accuracy.
[…]
Finding 3:  Longer CoT does not guarantee higher accuracy. Across all three models, failed generations consistently exhibit longer CoT lengths than successful ones (Figure 1). This indicates that excessively long reasoning chains often introduce redundancy and error propagation rather than improving performance, highlighting the need for appropriately controlled CoT lengths.
[…]
Long CoT introduces application latency. Excessively long CoT reasoning introduces significant latency overhead, which ultimately degrades user experience. To quantify this impact, we evaluate the inference performance of Deepseek-R1-Distill-Qwen models of varying sizes using the transformers library on a single NVIDIA A800 GPU. The measured generation speeds range from 46 tokens/s for the smallest (1.5B) model to 18 tokens/s for the largest (32B) model. Using these measurements, we estimate the average end-to-end latency on two benchmarks: HumanEval, representing relatively simple coding tasks, and Codeforces, which contains more complex algorithmic challenges. The results are presented in Figure 2.
Figure 2. Inference latency of different model sizes on HumanEval and Codeforces datasets with and without CoT reasoning. While CoT improves reasoning capability, it substantially increases application latency, particularly for larger models, posing practical constraints on real-world deployment.
Figure 2 shows that incorporating CoT substantially increases response time, particularly for larger models on complex tasks. On Codeforces, the 32B model with CoT requires an average of 776 seconds to complete the reasoning for each problem, compared to 157 seconds without CoT. Even the smaller 1.5B model takes more than five minutes per problem when using CoT. These delays make real-time use impractical. Moreover, all experiments were conducted under a 16k-token context limit, whereas Codeforces samples already exceed 14k tokens on average, implying that real-world latency would be even higher. Thus, while CoT improves reasoning, its verbosity imposes severe constraints on latency-sensitive applications, highlighting the need for effective CoT length control.
[…]
Finding 4:  Long CoT substantially slows down inference. For example, the 32B model on Codeforces requires nearly five times longer with CoT than without. This latency overhead makes CoT impractical for real-time or latency-sensitive applications, underscoring the need for effective length control.
4. Methodology
We propose SEER, a self-enhancing framework designed to enable LLMs to refine their reasoning capabilities by learning from their own generated outputs. Unlike conventional supervised fine-tuning pipelines that depend heavily on high-quality, human-labeled CoT annotations, SEER operates in a fully autonomous manner: the model generates candidate reasoning traces, filters them based on correctness and conciseness, and then learns from the selected high-quality examples. As illustrated in Figure 3, the framework is composed of three key stages: (1) Pre-inference generation of CoT responses, (2) BoN sampling for high-quality reasoning path selection, and (3) Adaptive CoT filtering to control verbosity without sacrificing reasoning fidelity. Through this design, SEER significantly reduces the reliance on prompt engineering and external compression tools, while promoting reasoning traces that are both accurate and concise, and that remain tailored to the nature of each target task.
Figure 3. An overview of the SEER framework, illustrating its self-enhancing procedure of generation, selection, and adaptive filtering.
4.1. Pre-Inference Data Generation
The first stage of SEER focuses on constructing a comprehensive corpus of CoT outputs to serve as the foundation for subsequent filtering and fine-tuning. In this step, the base language model is used to process each question in the training set, generating a complete solution that includes an explicit, step-by-step reasoning path.
A distinctive advantage of our approach is its low sensitivity to prompt engineering. While a prompt is required to elicit CoT reasoning, our subsequent adaptive filtering stage ensures that the method is not critically dependent on the choice of prompt format. This property allows us to adopt a pragmatic strategy: whenever an official, dataset-specific prompt template is available, we use it directly; otherwise, we revert to a robust general-purpose template. For example, for mathematical reasoning tasks, we utilize the system prompt from the study by Wei et al. (Wei et al., 2022):
[…]
To prevent incomplete reasoning caused by premature output truncation, we allocate a sufficiently high maximum token budget during generation. This ensures that the model’s natural CoT length distribution is faithfully captured. Accurately profiling this distribution is critical for calibrating the adaptive filter in later stages, as it allows us to set dataset-specific compression thresholds that minimize information loss.
4.2. Best-of-N Sampling for Data Refinement
To prevent the reasoning loops noted in Section 3, which can harm data quality, we adopt a BoN sampling strategy to build a high-quality training set. This strategy also helps obtain cleaner datasets with more concise CoT lengths. For each input question, the model generates $\mathcal{N}$ candidate responses, each augmented with a CoT. These candidates are then ranked and filtered according to a strict hierarchical set of criteria:
  1. (1)
    Correctness of the Final Answer: Only candidates producing the correct final answer are considered. All incorrect outputs are discarded.
  2. (2)
    Presence of Valid CoT: From the remaining candidates, we retain only those containing a non-empty reasoning path. This guarantees that each example is explainable and preserves its interpretability.
  3. (3)
    Conciseness of Reasoning: If multiple candidates satisfy the first two criteria, we select the one with the shortest CoT length. This prioritizes efficiency and prevents the inclusion of unnecessarily verbose reasoning.
This multi-stage filtering process ensures that the resulting dataset consists exclusively of reasoning traces that are (1) correct in outcome, (2) logically explicit, and (3) concise. By integrating BoN sampling into our pipeline, we are able to suppress reasoning loops, improve data quality, and provide the model with more effective fine-tuning signals.
4.3. Adaptive CoT Filtering
Even after BoN sampling, CoT responses can vary widely in length depending on the task, domain, and problem complexity. Excessively long CoTs waste computational resources, while overly aggressive compression risks discarding essential reasoning steps. To balance these factors, we introduce an adaptive CoT filtering mechanism that sets a dataset-specific maximum CoT length, denoted $\lambda_{c}$.
A fixed, manually chosen $\lambda_{c}$ is often suboptimal because optimal CoT lengths differ substantially across domains (e.g., mathematical proofs vs. code debugging). If $\lambda_{c}$ is too low, important intermediate reasoning may be lost, potentially reducing accuracy and even causing the model to overfit to overly compressed reasoning patterns. Conversely, a value set too high will dilute the benefits of compression, leading to minimal gains in efficiency.
Our adaptive filtering addresses this by first generating CoT responses for the entire training set during the pre-inference stage and then computing two statistical measures of their length distribution: the mean length $\bar{\lambda}$ and the median length $\tilde{\lambda}$. We define $\lambda_{ada}$ as:
$\lambda_{ada}=\frac{\bar{\lambda}+\tilde{\lambda}}{2}$
This formula anchors the threshold to the observed characteristics of each dataset, balancing robustness (median) with representativeness (mean).
While $\lambda_{c}$ provides a principled baseline, our design remains flexible. Users can manually adjust it to match specific efficiency or accuracy requirements. This tunable control allows practitioners to trade off between computational savings and reasoning completeness, enabling SEER to adapt seamlessly across diverse reasoning scenarios.
4.4. Overall Procedure
Input: Training set $\mathcal{D}=\{(x_{i},y_{i}^{\star})\}$, base model $\mathcal{M}$, candidates per input $\mathcal{N}$
Output: The model $\mathcal{M}_{\text{adapt}}$ fine-tuned on the filtered dataset $\widetilde{\mathcal{D}}$
Stage 1: Pre-inference generation;
foreach $(x_{i},y_{i}^{\star})\in\mathcal{D}$ do
  Generate $\mathcal{N}$ CoT-augmented candidates with $\mathcal{M}$ under a fixed prompt;
  $r_{ij}\leftarrow\text{final response generated for }x_{i}$;
  $\text{CoT}_{ij}\leftarrow\text{reasoning trace (Chain-of-Thought) leading to }r_{ij}$;
  $\mathcal{C}_{i}\leftarrow\{\,(r_{i1},\text{CoT}_{i1}),\ldots,(r_{i\mathcal{N}},\text{CoT}_{i\mathcal{N}})\,\}$;
Stage 2: Best-of-N (BoN) sampling;
foreach $(x_{i},y_{i}^{\star}),\mathcal{C}_{i}$ do
  $\mathcal{S}_{i}\leftarrow\{(r,\text{CoT})\in\mathcal{C}_{i}\mid\textsc{Correct}(r,y_{i}^{\star})\ \wedge\ |\text{CoT}|>0\}$ if $\mathcal{S}_{i}\neq\emptyset$ then
     $(r_{i}^{\star},\text{CoT}_{i}^{\star})\leftarrow\arg\min_{(r,\text{CoT})\in\mathcal{S}_{i}}|\text{CoT}|$ Add $(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})$ to $\mathcal{D}^{\prime}$;
Stage 3: Adaptive CoT filtering;
Compute length set $L\leftarrow\{\,|\text{CoT}_{i}^{\star}|\mid(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})\in\mathcal{D}^{\prime}\,\}$;
$\bar{\lambda}\leftarrow\text{mean}(L),\ \tilde{\lambda}\leftarrow\text{median}(L)$;  $\lambda_{c}\leftarrow(\bar{\lambda}+\tilde{\lambda})/2$;
$\widetilde{\mathcal{D}}\leftarrow\{(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})\in\mathcal{D}^{\prime}\mid|\text{CoT}_{i}^{\star}|\leq\lambda_{c}\}$;
Stage 4: Fine-tuning;
Fine-tune model on $\widetilde{\mathcal{D}}$ with SFT to obtain $\mathcal{M}_{\text{adapt}}$;
return $\mathcal{M}_{\text{adapt}}$;
Algorithm 1 Self-Enhancing Training Process with BoN & Adaptive CoT Filtering
Algorithm 1 summarizes the end-to-end process, from pre-inference generation to BoN-based selection, adaptive filtering, and final fine-tuning on the curated dataset.
Through the combination of high-coverage pre-inference generation, strict BoN filtering, and adaptive CoT length control, SEER produces training data that is both high-quality and efficiency-oriented. This not only improves the reasoning efficiency of the fine-tuned model but also mitigates failure modes such as infinite reasoning loops and truncation.
5. Experimental Setup
5.1. Datasets
To thoroughly evaluate the performance and efficacy of our method, we conduct experiments across three distinct software engineering tasks and one mathematical reasoning task. These tasks were selected to cover a broad spectrum of problem types, input–output modalities, and reasoning requirements. By incorporating both code-related and mathematical reasoning benchmarks, we aim to investigate not only the in-domain effectiveness of our method but also its adaptability to different reasoning paradigms and task formats.
  • MathQA-Python (Austin et al., 2021) is a dataset designed for generating Python programs that solve mathematically formulated problems. Each problem is presented in natural language and requires the model to produce an executable Python solution. The challenges in this dataset include parsing problem statements, performing intermediate symbolic reasoning, and implementing the reasoning steps as correct Python code. This dataset serves as a baseline for evaluating code generation in mathematically oriented problem-solving contexts.
  • CodeXGLUE-Defect-Detection (Zhou et al., 2019) is a binary classification dataset that assesses a model’s ability to identify whether a given piece of source code contains security vulnerabilities or functional defects. The data originates from real-world code bases, and the detection task often requires understanding complex control flows, data dependencies, and security implications. This dataset evaluates the model’s capability in static code analysis and secure software engineering scenarios, where concise yet accurate reasoning is critical to avoid excessive inference overhead.
  • Code-Search-WebQuery (Lu et al., 2021; Husain et al., 2019) is used to test semantic code retrieval capabilities. In this task, the model is provided with a natural language query and a code snippet, and must determine whether the code correctly satisfies the described requirement. This problem is framed as binary classification, and it challenges the model to bridge the semantic gap between natural language descriptions and programming logic. The task reflects realistic software development workflows, such as searching for relevant functions in large code bases.
  • GSM-8K (Cobbe et al., 2021) is a large-scale benchmark for multi-step mathematical reasoning. Each question requires systematic decomposition into intermediate steps before arriving at the final answer. Unlike MathQA-Python, the output here is not executable code but a final numeric answer, making it an ideal testbed for evaluating whether our method can compress reasoning chains without degrading accuracy in purely symbolic reasoning scenarios. It also provides an out-of-domain evaluation for models fine-tuned on code-related tasks.
Additionally, in RQ3 we include the HumanEval dataset introduced in Section 3 to evaluate the generalizability of our method. By selecting this combination of datasets, our evaluation covers tasks ranging from code synthesis, defect detection, and semantic code retrieval to mathematical reasoning.
5.2. Implementation Details
All experiments are conducted on a server equipped with four NVIDIA RTX A800 GPUs (80 GB each) and 128 Intel Xeon CPUs, running Ubuntu 22.04 LTS. We fine-tune DeepSeek-R1-Distill-Qwen-7B model (DeepSeek-AI, 2025b) under a full-parameter supervised fine-tuning (SFT) setting, using the LLaMAFactory framework (Zheng et al., 2024). Training is performed for three epochs with an effective batch size of eight, gradient checkpointing enabled, and the AdamW optimizer. The learning rate is fixed at $1\times 10^{-5}$ with a cosine decay schedule and 10% warm-up, and gradient clipping (norm 1.0) is applied for stability. During fine-tuning, inputs are tokenized with a maximum sequence length of 8,192 tokens, truncating from the input side if necessary to preserve the output reasoning chain. Although we adopt full-parameter SFT by default, the SEER framework is also compatible with parameter-efficient fine-tuning methods such as LoRA and QLoRA, which we discuss in the Discussion section.
5.3. Baselines
In our work, we consider the following baselines.
  • Self-Training: Directly using self-inference data to train the model.
  • TokenSkip (Xia et al., 2025): Fine-tuning LLMs by selectively skipping less important tokens while retaining key tokens, enabling the generation of compressed CoTs with adjustable ratios.
  • Naive BoN (Munkhbat et al., 2025b): Use only the Best-of-N method to process the pre-inference data and then train the model.
In order to ensure the fairness of the experiment, we have implemented several modifications to the TokenSkip source code. These changes include: 1) the incorporation of the “deepseek” prompt template, as this template is employed by the chosen model; and 2) the use of the vllm library in place of the transformers library to achieve more efficient inference and to mitigate any potential precision loss that could lead to unfair comparisons.
5.4. Evaluation Metrics
We measure the following metrics:
[…]
Accuracy: We use the same pass@1, which is mentioned in Section 3 as the primary metric to evaluate correctness. For programming tasks (e.g., MathQA-Python), a prediction is considered correct if the generated code passes all test cases. For mathematical reasoning tasks (e.g., GSM-8K), we adopt exact matching criterion, where the model’s final answer must exactly match the ground truth.
[…]
Compression Rate: We define the Compression Rate $\rho$ to quantify the conciseness of the generated CoT. Specifically, $\rho$ measures the relative reduction in the number of tokens compared to a reference CoT. It is computed as: $\rho=(\lambda-{\lambda^{\prime}})/{\lambda}$
In this formula, $\lambda$ denotes the token count of the reference (original) CoT, and $\lambda^{\prime}$ represents the token count of the CoT generated by the evaluated model or baseline. A higher value of $\rho$ indicates a greater degree of compression, corresponding to a more concise reasoning path. This metric is particularly important for evaluating methods such as TokenSkip and Naive BoN, which are explicitly designed to produce shorter CoTs without compromising correctness.
6. Experiment
In this section, we conduct experiments to answer the following research questions.
6.1. Research Questions
In order to evaluate our work, we address the following research questions:
  • RQ1: How does the CoT compression performance of SEER compare with baselines? We evaluate SEER against baseline methods on multiple datasets.
  • RQ2: What is the contribution of each component of SEER? We conduct ablation studies by removing or varying parts of SEER, such as the CoT length filter $\lambda_{c}$ and the BoN parameter $\mathcal{N}$, to examine their impact.
  • RQ3: How well does SEER generalize across domains? We examine whether SEER remains effective when the fine-tuned data and test data come from different domains.
  • RQ4: Can SEER mitigate reasoning loops? We test whether SEER helps LLMs generate valid answers without falling into infinite reasoning loops.
6.2. RQ1: The Performance of SEER
In this subsection, we assess the performance of SEER in comparison with baseline methods. For the TokenSkip experiment, we set $\gamma=0.5$, meaning that we append the sequence “[EOS]0.5[EOS][EOS]0.5[EOS]” to the end of the input prompt. Similarly, for Naive BoN, we set $N=3$, which is consistent with the parameters used in evaluating our method. Table 2 reports task pass@1 accuracy ($\mathcal{A}$), average CoT length ($\lambda_{\text{avg}}$), and compression rate ($\rho$) under the 8K token constraint.
Methods MathQA-Python Defect-Detection Code-Search GSM-8K $\rho_{avg}(\%)$
$\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$
Base 63.7 1,456 - 44.7 1,836 - 72.4 472 - 90.1 1,309 - -
Self-Trained 74.6 1,177 19.2 49.3 1,055 42.5 72.5 468 0.9 89.2 792 39.5 25.5
Naive BoN (N=3) 73.6 1,191 18.2 49.8 1,042 43.3 74.9 478 -1.3 89.8 845 35.5 23.9
TokenSkip ($\gamma$=0.5) 1.6 1,196 17.9 36.2 2,421 -31.9 66.5 465 1.5 37.9 1,238 5.4 -1.8
TokenSkip ($\gamma$=0.8) 0.9 1,109 23.8 47.2 1,004 45.3 76.4 341 27.8 87.3 1,197 8.6 26.4
SEER 74.9 877 39.8 50.5 785 57.2 77.3 362 23.3 90.5 677 48.3 42.1
    The Base model refers to “DeepSeek-R1-Distill-Qwen-7B”. The same applies to the other tables;
    $\mathcal{A}(\%)$: Pass@1; $\lambda_{avg}$: Average output length (tokens);
    $\rho(\%)$: Compression rate compared to the Base model; $\rho_{avg}(\%)$: Average compression rate across all datasets.
Table 2. The CoT Compression Performance Over Different Datasets (8K Token Limit)
Across all three software engineering tasks, SEER shows strong performance in both accuracy and CoT compression. On MathQA-Python, which requires code generation, it achieves 74.9% pass@1 with a 39.8% reduction in reasoning length. On Defect-Detection, it obtains the best pass@1 (50.5%) and the highest compression rate (57.2%). On Code-Search, it maintains a high pass@1 of 77.3% while compressing CoTs by 22.9%. Beyond software engineering, SEER also performs well on GSM-8K, a math reasoning dataset, reaching 90.5% pass@1 with a 48.3% reduction in CoT length. The accuracy gains are largely due to the mitigation of truncation and excessive reasoning, which often undermine model reliability. These results demonstrate that SEER consistently balances brevity and accuracy across diverse tasks.
[…]
Compared to TokenSkip, both Self-Trained and Naive BoN achieve more stable outcomes but still fall short of SEER. On Defect-Detection, their pass@1 remains below 50%, with compression rates under 30%. While they provide modest improvements over the base model, they cannot match SEER’s combination of high accuracy and substantial compression. This suggests that generic fine-tuning or simple BoN sampling is insufficient for the challenging trade-off between performance and efficiency.
[…]
Answer to RQ1: Compared with all baselines, SEER consistently achieves high accuracy and efficient CoT compression on both software engineering tasks and the math reasoning task. On average, SEER reduces CoT length by 42.1%, representing a 16.9% improvement over the best existing baseline, while preserving or even improving accuracy across diverse tasks.
6.3. RQ2: The Impact of Each Components
To investigate the contribution of each component, we conduct ablation studies on the MathQA-Python dataset by varying the BoN sampling size $\mathcal{N}$ and the CoT length filter threshold $\lambda_{c}$. The results are summarized in Table 3.
Table 3. The Ablation Results of SEER on MathQA-Python Dataset
Methods Settings Pass@1 $\mathcal{A}$ (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
Base - 63.67 1,456 -
Without Filter $\mathcal{N}=1$ 73.49 1,191 18.20
$\mathcal{N}=3$ 74.45 1,177 19.16
$\mathcal{N}=5$ 74.40 1,163 20.12
Without BoN $\lambda_{c}=2\times\lambda_{ada}$ 74.40 1,071 26.44
$\lambda_{c}=1.5\times\lambda_{ada}$ 74.19 918 36.95
$\lambda_{c}=\lambda_{ada}$ 73.50 878 39.70
$\lambda_{c}=0.5\times\lambda_{ada}$ 72.76 836 42.58
SEER
$\mathcal{N}=3$ $\lambda_{c}=\lambda_{ada}$
74.88 877 39.77
When only BoN sampling is applied without the filter, increasing $\mathcal{N}$ from 1 to 3 improves accuracy from 73.49% to 74.45%, while the average CoT length decreases slightly from 1,191 to 1,177 tokens. However, further increasing $\mathcal{N}$ to 5 does not provide additional benefit and even slightly reduces accuracy (74.40%), indicating that the advantage of BoN sampling quickly saturates. In this setting, the compression rate remains relatively modest, ranging from 18.20% to 20.12%.
In contrast, applying only the adaptive filter leads to a substantial increase in compression. For instance, when $\lambda_{c}=2\times\lambda_{ada}$, the compression rate reaches 26.44%, and as the threshold becomes stricter ($\lambda_{c}=0.5\times\lambda_{ada}$), compression improves further to 42.58%. Nevertheless, overly aggressive filtering reduces accuracy, dropping from 74.40% to 72.76%. These results show that while the filter is more effective than BoN in reducing the redundancy in CoT, it risks discarding essential reasoning steps when applied too strongly.
To balance accuracy and compression rate, we set $\mathcal{N}=3$ and $\lambda_{c}=\lambda_{ada}$ in our method. When both components are combined in SEER with these settings, the method achieves the best trade-off, with the highest accuracy of 74.88% and a competitive compression rate of 39.77%. This demonstrates that the two modules are complementary: BoN sampling enhances robustness, while the adaptive filter enforces compactness, and together they enable SEER to deliver concise yet accurate reasoning.
[…]
Answer to RQ2: The ablation results reveal that BoN sampling alone offers only limited gains, while the adaptive filter strongly enhances compression but may harm accuracy if too strict. Their combination achieves the best trade-off, yielding both higher accuracy and effective compression, highlighting the complementary roles of the two components.
6.4. RQ3: Generalizability of SEER
In this section, we evaluate the generalizability of SEER by fine-tuning it on one dataset and testing on the remaining three datasets. The results are presented in Table 4.
Table 4. The Generalizability Evaluation Result of SEER Across All Tasks
MathQA-Python Defect-Detection Code-Search GSM-8K
Models $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$
Base 63.7 1456 - 44.7 1,836 - 72.4 472 - 90.1 1,309 -
MathQA-Py Fine-tune 74.9 877 39.8 47.1 1131 38.4 77.3 395 16.3 87.2 748 42.9
Defect-Detection Fine-tune 64.2 1279 12.2 50.5 785 57.2 72.2 339 28.2 90.4 1128 13.8
Code-Search Fine-tune 64.1 1104 24.2 51.4 672 63.4 77.3 362 23.3 89.8 853 34.8
GSM-8K Fine-tune 62.5 1138 21.8 48.5 1483 19.2 79.6 435 7.8 90.5 677 48.3
Overall, the findings demonstrate that SEER exhibits strong cross-task generalization. Regardless of which dataset is used for fine-tuning, the resulting models generally achieve higher accuracy compared with the base model. Only a few cases show a slight decrease, such as fine-tuning on GSM-8K and then evaluating on MathQA-Python (62.5% vs. 63.7%). In all cases, the compression rate remains consistently above zero. For example, fine-tuning on MathQA-Python compresses average reasoning length on GSM-8K from 1,309 tokens to 748 tokens (42.9% compression), while still preserving high accuracy (87.2%). This indicates that SEER is able to reduce reasoning length while maintaining competitive accuracy even on out-of-domain tasks.
Interestingly, in some settings, the fine-tuned models even surpass those trained directly on their target dataset. For instance, the model fine-tuned on Code-Search achieves 51.4% accuracy on Defect-Detection, which is higher than the model fine-tuned on Defect-Detection (50.5%). These results suggest that SEER captures reusable reasoning patterns that extend beyond the dataset it was trained on, rather than overfitting to a single task.
Figure 4. Performance of SEER fine-tuned on different datasets when evaluated on HumanEval. Fine-tuning on coding-related datasets such as MathQA-Python and Code-Search yields the best generalization with shorter CoT lengths, while non-coding datasets like GSM-8K still provide moderate improvements over the base model.
To further assess its generalizability, we evaluate the four fine-tuned models on the HumanEval dataset, as shown in Figure 4. The results reveal two key insights. First, all fine-tuned models achieve noticeable improvements in both accuracy and CoT length over the base model, confirming that the knowledge learned from different tasks generalizes to program synthesis. Second, the reduction in reasoning length remains consistent, suggesting that SEER’s compression mechanism is stable even when applied to tasks outside the training domain. This demonstrates that the framework does not merely adapt to a specific dataset but instead learns reasoning strategies that generalize across diverse software engineering and mathematical tasks.
[…]
Answer to RQ3: SEER demonstrates strong cross-task generalization. Even when evaluated on out-of-domain datasets, it consistently improves accuracy while reducing CoT length. Fine-tuning on coding datasets further generalizes well to HumanEval, confirming that SEER learns reusable reasoning strategies beyond its training domain.
6.5. RQ4: The Ability of Reducing Reasoning Loops
Infinite reasoning loops occur when a model repeatedly generates identical reasoning segments until the output reaches the maximum context length and is truncated. This phenomenon wastes computational resources, inflates latency, and often produces incomplete or invalid final answers. Because such loops can arise unpredictably across models of different sizes, they represent a critical robustness concern. In this section, we investigate whether our approach can mitigate this problem.
For each dataset, we adapt our method to fine-tune a model specific to that dataset and evaluate it against the base model. The results in Table 5 show that SEER substantially reduces reasoning loops in all cases. On MathQA-Python, loop occurrences drop from 85 to 23 (72.9% reduction). On Code-Search, the count falls from 15 to 1 (93.3% reduction), effectively eliminating loops. The largest improvement appears in Defect-Detection, where loops decrease from 222 to just 7 (96.8% reduction). Even on GSM8K, which generally contains shorter reasoning chains, loops drop from 54 to 6 (88.9% reduction).
Table 5. The Performance of SEER in Reducing Truncation and Reasoning Loops
MathQA-Python Code-Search Defect-Detection GSM8K
Truncation Loops Truncation Loops Truncation Loops Truncation Loops
Base 110 85 18 15 235 222 59 54
SEER 30 23 1 1 7 7 8 6
During our empirical analysis in Section 3, we find that most truncations in the base model are directly caused by such loops. This behavior wastes much of the available context budget and prevents the model from completing its reasoning. Our adaptive filtering strategy compresses CoT length, which mitigates loop formation and reduces truncation, as shorter reasoning provides fewer chances for repetitive output. This dual effect, loop suppression and truncation reduction, further enhances the robustness and practicality of our method in real-world deployments.
Overall, these results confirm that SEER offers a robust defense against infinite reasoning loops, complementing its primary goal of efficient CoT compression and providing benefits for latency, reliability, and output quality.
[…]
Answer to RQ4: SEER effectively suppresses infinite reasoning loops across all datasets, with reductions ranging from 72.9% on MathQA-Python to 96.8% on Defect-Detection. Because most truncations in the base model stem from such loops, the adaptive filtering strategy not only prevents pathological repetition but also lowers truncation frequency.
7. Discussion
7.1. The Impact of Training Approaches
An important aspect of our framework is its adaptability to different fine-tuning paradigms. Beyond full-parameter supervised fine-tuning (SFT), we also explored parameter-efficient alternatives such as LoRA. The comparison reveals two insights. First, while SFT unsurprisingly achieves the strongest overall performance, LoRA retains most of the benefits of our method, delivering substantial accuracy gains and compression improvements over the base model. This indicates that the effectiveness of adaptive CoT filtering does not depend on updating the entire parameter set.
Table 6. The Performance Comparison Between Using SFT and PEFT Fine-tuning Method
Pass@1 $\mathcal{A}$ (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
Base 63.67 1,456 -
PEFT (LoRA) 67.71 949 34.8
SFT 74.88 877 39.8
Second, the LoRA-based approach demonstrates that our approach remains effective in environments with limited resources. By reducing computational and storage costs, LoRA enables deployment in scenarios where full SFT would be prohibitively expensive. Although the absolute gains are somewhat smaller than those under SFT, the efficiency–effectiveness trade-off makes PEFT methods an attractive option for real-world applications.
Taken together, these observations indicate that our framework is effective under ideal training conditions and remains flexible in limited-resource settings. This adaptability broadens its applicability, making it feasible to integrate SEER into a wider range of deployment scenarios.
7.2. Threats to Validity
7.2.1. Internal Validity
Internal validity concerns whether the observed improvements of SEER truly arise from our proposed framework rather than confounding factors. One potential threat is the reliance on a single model family, namely the DeepSeek-R1-Distill-Qwen-7B. Although we carefully fine-tuned and evaluated under consistent settings to reduce bias, the conclusions may partially reflect properties of this architecture rather than general effectiveness. To mitigate this risk, we performed ablation studies (varying BoN sampling and adaptive filtering thresholds) to confirm that the observed gains stem from the specific design choices of SEER rather than incidental artifacts. Future work will extend the evaluation to other model families, including implicit CoT models that do not explicitly output reasoning traces. Moreover, testing on closed-source models with advanced reasoning capabilities (e.g., OpenAI O1, Gemini, Claude’s Thinking series) will help validate whether the improvements are architecture-independent.
7.2.2. External Validity
External validity concerns the generalizability of our findings beyond the datasets and settings considered in this work. Our experiments primarily focus on software engineering tasks, as they represent one of the most critical application domains for reasoning models. To mitigate the risk of over-specialization, we additionally included a mathematical reasoning dataset (GSM-8K), which provides a different reasoning style and serves as an initial step toward broader validation. Nevertheless, important domains such as commonsense reasoning, legal QA, and scientific problem solving remain unexplored.
Moreover, all evaluations were conducted with fixed prompts and prompting strategies. While our method is designed to be prompt-agnostic, alternative instruction formats, few-shot configurations, or domain-specific prompts could influence the observed trade-off between compression and accuracy. Finally, although the datasets used contain tasks of varying difficulty, we have not systematically examined whether compression benefits differ across problem complexity levels.
To further strengthen external validity, we plan to expand our evaluation to additional reasoning domains and more diverse datasets, as well as investigate robustness under alternative prompting strategies. Future work will also include stratified analyses by task difficulty to better understand whether SEER is equally effective on both simple and challenging instances.
8. Related Works
Since the concept of CoT was first introduced (Wei et al., 2022), research on controlling and optimizing CoT length has drawn increasing attention. Numerous studies have shown that extending intermediate reasoning steps within CoT prompting can significantly enhance the capabilities of LLMs (Fu et al., 2022; Merrill and Sabharwal, 2023; Jin et al., 2024). However, other work (Wu et al., 2025) indicates that performance does not increase indefinitely; instead, it follows a rise-and-fall pattern as CoT length grows. This suggests that optimally controlling CoT length is a key challenge for balancing reasoning quality and efficiency.
$\blacksquare$ Prompt-Guided Efficient Reasoning. Prompt-based approaches improve reasoning efficiency by explicitly guiding the model to be concise. CCoT (Cheng and Van Durme, 2024) uses a concise CoT template to encourage brevity through prompt wording. Token-Budget (Han et al., 2024) proposes a token-budget-aware reasoning framework that dynamically adjusts reasoning length based on estimated complexity. CoD (Xu et al., 2025a) limits reasoning to within five words via a “Chain-of-Draft” template. Token Complexity (Lee et al., 2025) systematically analyzes prompt compression strategies, highlighting that many still operate far from theoretical efficiency limits.
These methods can reduce verbosity without modifying model parameters, but they are inherently sensitive to prompt design and task distribution. In contrast, our method removes this dependency: by learning from filtered, concise, and correct CoT examples, the model internalizes concise reasoning patterns that generalize across prompts and domains.
$\blacksquare$ Model-Optimization Based Efficient Reasoning. Other approaches modify the model or its training data to achieve concise reasoning. C3oT (Kang et al., 2025) trains models jointly on long and short CoTs to learn mappings between them, using conditioned inference to produce short CoTs. TokenSkip (Xia et al., 2025) enables skipping less important tokens during generation, allowing controllable compression but sometimes causing instability (e.g., reasoning loops in our observations). LightThinker (Zhang et al., 2025) compresses verbose thoughts into compact internal representations, discarding the original CoT from the context. CODI (Shen et al., 2025) distills CoT into a continuous space, jointly learning explicit and implicit reasoning traces.
While these methods reduce reasoning length, many require additional compression modules, complex training objectives, or implicit representations that limit interpretability. Our method is simpler: it uses the model’s own generated data, combined with a lightweight Best-of-N sampling and adaptive filtering, to directly fine-tune for concise yet explicit reasoning. This avoids extra components and preserves the transparency of reasoning steps.
$\blacksquare$ Chain-of-Thought Effectiveness in Software Engineering. CoT has shown strong benefits in software engineering tasks. SCoT (Li et al., 2025) was among the first to apply structured CoT to code generation. The emergence of Deepseek-R1 (DeepSeek-AI, 2025b) further highlighted the value of streamlined CoT for software engineering tasks. Liu et al. (Liu et al., 2025) revisited CoT training paradigms, showing that some seemingly effective methods do not transfer well to real-world coding tasks. UnCert-CoT (Zhu et al., 2025) uses uncertainty estimation to decide when to trigger CoT, avoiding unnecessary reasoning. CoD (Yang, 2025) adapts the “Chain-of-Draft” approach to software engineering tasks, greatly reducing token usage.
However, existing SE-focused methods often target specific task scenarios or rely on inference-time control, which may not eliminate pathological failure modes, such as infinite reasoning loops. Our method differs by explicitly filtering out redundant or looping reasoning during training, allowing the model to internalize concise, stable reasoning behaviors. This makes it robust not only in software engineering tasks but also in mathematical reasoning, while retaining interpretability and avoiding the need for task-specific prompt engineering.
9. Conclusion
Our work systematically examined the inefficiency of excessively long CoT reasoning, identifying that long CoT can lead to truncation and overthinking, resulting in decreased accuracy. We then present SEER, a self-enhancing framework for adaptive CoT compression. Through extensive experiments on software engineering and mathematical reasoning tasks, SEER reduces CoT length by an average of 42.1% while maintaining or even improving accuracy. Our ablation results confirm the complementary roles of BoN sampling and adaptive filtering, and cross-domain evaluations demonstrate SEER’s strong generalizability. SEER also effectively mitigates truncation and suppresses infinite reasoning loops, addressing key efficiency and robustness issues. These findings highlight SEER as a practical approach for making CoT-enhanced LLMs more efficient and reliable. Future work will extend SEER to broader reasoning domains, apply it to diverse model architectures, and explore integration into latency-sensitive real-world systems.

### v2

@@
 Abstract.
-The introduction of Chain-of-Thought (CoT) reasoning has significantly enhanced the problem-solving capabilities of Large Language Models (LLMs) by prompting them to articulate intermediate reasoning steps rather than producing answers directly.
-This structured approach has led to significant improvements in accuracy and robustness across benchmarks in arithmetic, logic, and commonsense reasoning.
-Although effective, it introduces computational overhead.
-Each reasoning step requires additional token generation, which raises inference latency, memory usage, and key-value cache demands.
-These issues are particularly critical in domains such as software engineering, where deterministic logic and concise outputs are preferred.
-To understand how CoT length impacts both model performance and user experience, we conduct an empirical study on widely used code generation datasets.
-We find that longer CoT does not always improve performance.
-Across benchmarks, excessively long reasoning often results in truncation, reduced accuracy, and latency increases of up to five times, with failed outputs consistently longer than successful ones.
-These findings challenge the assumption that longer reasoning is inherently beneficial and underscore the need for adaptive CoT control.
-Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive and self-enhancing framework for CoT compression that improves the conciseness of reasoning while preserving output quality, without relying on external compression tools.
-SEER follows an adaptive self-enhancing paradigm, where CoT responses are refined by combining Best-of-N sampling with task-aware adaptive filtering.
-The adaptive filter dynamically adjusts thresholds based on the token length of pre-inference model outputs, effectively controlling verbosity while reducing computational costs and preserving accuracy.
-We evaluate SEER on three software engineering tasks and one mathematical reasoning task.
-On average, SEER reduces CoT length by 42.1% without compromising performance, and accuracy even improves due to reduced truncation.
-Moreover, it substantially lowers truncation rates and eliminates most infinite reasoning loops, thereby enhancing both efficiency and robustness.
-These results highlight SEER as a practical approach to making the use of CoT more efficient, enabling CoT-enhanced LLMs to operate effectively even in resource-constrained settings and on complex tasks.
-Chain-of-Thought (CoT), Large Language Models (LLMs), CoT Compression, Software Engineering, Reasoning Efficiency
+Chain-of-Thought (CoT) prompting can substantially improve the reasoning ability of large language models (LLMs), but it often comes with high inference cost due to long and poorly controlled reasoning traces.
+This overhead is particularly problematic in software engineering tasks (e.g., code generation), where both latency and output reliability matter.
+To better understand this trade-off, we conduct an empirical study on widely used code generation benchmarks and observe that many modern reasoning models produce excessively verbose CoTs (often thousands of tokens), which frequently leads to truncation and unstable generation.
+Using a strict $n$-gram repetition detector, we find that the vast majority of truncations are associated with degenerate looping behaviors.
+Moreover, longer CoT does not necessarily yield better outcomes: failed generations tend to be longer than successful ones, indicating diminishing or even negative returns from overlong reasoning.
+Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), a self-enhancing framework for adaptive CoT compression.
+SEER improves the conciseness of reasoning while preserving output quality, without relying on external compression tools.
+SEER refines self-generated CoT data via Best-of-N sampling to suppress looping and redundant traces, then applies a lightweight, data-driven filter to encourage concise yet correct reasoning.
+It then fine-tunes the model on the filtered data to internalize concise reasoning behaviors.
+Across three software engineering tasks, SEER reduces CoT length by 41.6% on average while improving task performance, largely by reducing truncation and mitigating reasoning loops.
+Chain-of-Thought Compression, Software Engineering, Large Language Models, Reasoning Efficiency
 copyright: none
@@
 The introduction of Chain-of-Thought (CoT) marks a pivotal advancement in enhancing the reasoning capabilities of Large Language Models (LLMs), particularly for complex, multi-step tasks.
-By encouraging models to explicitly articulate intermediate reasoning steps, CoT enables a systematic, step-by-step breakdown of problems rather than merely producing an answer in a single attempt (Kojima et al., 2022; Wei et al., 2022).
-While CoT has demonstrated notable improvements in accuracy and robustness across tasks involving arithmetic, logic, and commonsense reasoning (Nye et al., 2021; Wang et al., 2023; Zhou et al., 2022), its effectiveness becomes even more pronounced in software engineering tasks.
-Such tasks typically demand precise logical reasoning, structured multi-step problem solving, and transparent decision-making, all of which are well supported by CoT.
-Recent work shows that CoT can substantially improve performance in multiple software engineering tasks (e.g., code generation, program synthesis, and automated debugging) by improving the interpretability of the model’s reasoning and aligning it with human problem-solving processes (Yao et al., 2023; Li et al., 2025).
-However, this benefit comes with substantial costs.
-As most contemporary LLMs are based on the Transformer architecture (Vaswani et al., 2017), the attention mechanism incurs quadratic computational complexity with respect to context length, while the KV cache grows linearly with context size.
-Consequently, the verbose decoding of CoT significantly increases inference latency and memory usage, with KV-cache requirements emerging as a key bottleneck (Agrawal et al., 2024; Ho et al., 2024).
-This inefficiency is particularly critical in software engineering tasks, where timely responses are essential for developer productivity and effective decision-making.
-Moreover, standard CoT often “overthinks” even simple code generation tasks, generating excessive reasoning tokens despite the existence of straightforward solutions.
-In more complex scenarios (e.g., repairing software vulnerabilities), this overthinking can impair performance as it increases the inference costs and latencies (Cuadron et al., 2025).
-Consequently, reducing the length of CoT reasoning without compromising model performance remains a key challenge.
-Recent research indicates that compressing the CoT in training data and subsequently fine-tuning the model can effectively manage the length of the generated output while still maintaining accuracy (Kang et al., 2025; Munkhbat et al., 2025a; Xia et al., 2025).
-Methods such as TokenSkip allow LLMs to eliminate semantically redundant tokens, achieving significant reductions (e.g., from 313 to 181 tokens on the GSM-8K dataset) with minimal performance loss (Xia et al., 2025).
-Advanced frameworks like Adaptive GoGI-Skip, R1-Compress, and CoLaR have further improved these gains by reducing CoT lengths by 45%–80% while preserving accuracy, as well as delivering 1.6 to 2 times speedups in inference (Zhuang et al., 2025; Wang et al., 2025b; Tan et al., 2025).
-However, most of these methods are evaluated on mathematical reasoning datasets and commonsense QA, without assessing their effectiveness on software engineering problems; consequently, they may achieve good performance on mathematical reasoning tasks but exhibit substantially weaker results in software engineering contexts.
-Moreover, most of these approaches use external compression tools to compress CoT, such as GPT-4 in C3oT (Kang et al., 2025) and LLMLingua in TokenSkip (Xia et al., 2025), which may lead to information loss.
-Such information loss may lead to “thought leap” and ultimately influence the performance of the model (Xu et al., 2025b).
-For software engineering tasks, these compression tools may change the logic of the code or even make the code unable to execute.
-In this work, we conduct an empirical study to investigate the impact of excessively long CoT outputs on coding tasks using four open-source reasoning models.
-Our empirical study shows that excessively long CoT reasoning can harm the performance of LLMs.
-On HumanEval, small LLMs suffer substantial accuracy drops from truncation (e.g., 60.36% to 46.34% on the 1.5B model), while larger models show limited gains since failed outputs are often linked to longer reasoning.
-In addition, lengthy CoT raises latency by up to five times, underscoring the need to control reasoning length.
-Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive and self-enhancing framework for CoT compression that improves the conciseness of reasoning while preserving output quality, without relying on external compression tools.
-SEER adopts a self-enhancing paradigm, in which the model generates its own CoT-augmented responses and refines them via a two-stage data refinement process: 1) Best-of-N (BoN) sampling, which selects the shortest correct reasoning trace among multiple candidates to ensure both correctness and conciseness, and 2) an adaptive CoT filtering mechanism that dynamically calibrates maximum reasoning length based on the statistical properties of each dataset.
-This design enables task-specific control over verbosity, mitigating over-compression risks while eliminating pathological repetition patterns.
-Notably, SEER fine-tuned for one domain (e.g., code generation) also demonstrates strong transferability, preserving concise reasoning in other domains without retraining.
-We evaluate SEER on three diverse software engineering tasks and one mathematical reasoning benchmark under an 8K token context constraint.
-Experimental results show that SEER achieves an average CoT length reduction of 42.1% while preserving or improving accuracy across all tasks.
-Compared to baselines such as TokenSkip and Naive BoN, our method remains robust in complex software engineering tasks, where other approaches often fail due to instability or logical corruption.
-Moreover, SEER mitigates infinite reasoning loops, which occur when a model repeatedly generates the same fragment until the context limit is reached, by up to 97.7%, substantially reducing truncation rates.
+By encouraging models to explicitly articulate intermediate reasoning steps, CoT enables a step-by-step breakdown of problems rather than producing an answer in a single attempt (Kojima et al., 2022; Wei et al., 2022).
+Existing studies have shown that CoT achieves notable improvements across tasks, including arithmetic, logic, and commonsense reasoning (Nye et al., 2021; Wang et al., 2023; Zhou et al., 2022).
+Recently, the utilization of CoT has become more important in software engineering tasks, since they typically require precise logical reasoning, structured multi-step problem solving, and decision-making (Yao et al., 2023; Li et al., 2025).
+Figure 1. Examples of Common Chain-of-Thought Issues Across Different Models
+Figure 1 presents examples of common CoT issues observed across different open-source reasoning models.
+Even state-of-the-art reasoning LLMs may generate excessive or poorly controlled CoT, including overthinking, repetitive reasoning, and semantically unproductive deliberation.
+This issue can arise even in simple tasks with concise solutions.
+Such verbosity increases inference cost and latency, which is particularly problematic in code‑related scenarios.
+In software‑engineering contexts where reasoning is part of agent‑based workflows, uncontrolled CoT can increase system overhead and cause unnecessary reasoning steps, as observed in AutoGPT and LangGraph (Significant Gravitas, 2023a; LangChain AI, 2024; Significant Gravitas, 2023b).
+A common case is that an agent repeatedly re‑examines the same action or planning step without advancing, which wastes tokens and slows execution.
+To examine the effect of verbose CoTs in real-world settings, we conduct a systematic empirical analysis of CoT reasoning across a diverse set of open-source models on code generation tasks.
+We observe that excessively long CoT is common: most reasoning models generate 2K to 4K CoT tokens on average, and truncation can reach up to 17.1% under a 16K token budget.
+This verbosity is closely associated with reduced efficiency and reliability, as enabling CoT can increase token cost by up to an order of magnitude.
+Using a strict $n$-gram repetition detector, we find that 236 out of 261 truncations (90.4%) are identified as looping, and manual inspection suggests that this figure is a conservative lower bound.
+More importantly, our empirical results show that prompt-based conciseness control is limited and highly model-dependent.
+Overall, these findings highlight the need for effective mechanisms to regulate reasoning length and mitigate looping behaviors, i.e., CoT compression methods.
+Prior work on CoT compression can be generally divided into two categories, one of which applies explicit compression operations to CoT, for example, TokenSkip with LLMLingua and C3oT with GPT-4 (Xia et al., 2025; Kang et al., 2025; Zhuang et al., 2025).
+These methods usually achieve high compression ratios, but they often cause information loss and “thought leap” effects (Xu et al., 2025a).
+The second line of work relies on repeated sampling or distillation to collect simpler reasoning examples and encourage the model to produce more concise reasoning, such as Naive BoN (Munkhbat et al., 2025; Yu et al., 2024).
+This form of implicit compression avoids information loss, but typically cannot achieve a high level of compression.
+In this paper, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive framework designed to make CoT reasoning both effective and efficient for software engineering tasks.
+Rather than relying on external compression tools, SEER internalizes CoT control into the training process by learning concise reasoning patterns directly from self-generated outputs.
+SEER adopts a self-enhancing paradigm in which the model generates multiple CoT-augmented candidates and refines them through two complementary mechanisms.
+First, Best-of-N (BoN) sampling explicitly addresses looping behaviors by selecting the shortest correct reasoning trace among multiple candidates, effectively suppressing loops and redundant expansions observed in our empirical study.
+Second, SEER introduces an adaptive CoT filtering strategy that calibrates maximum reasoning length based on dataset-specific length distributions.
+This mechanism is grounded in the empirical observation that effective reasoning converges to a narrow length range across models, and prevents both over-compression and excessive verbosity.
+We evaluate SEER on three representative software engineering tasks: code generation, defect detection and natural language code search.
+Across all tasks, SEER reduces CoT length by an average of 41.6% while preserving or even improving pass@1 accuracy.
+Compared with existing compression baselines, SEER demonstrates superior robustness, achieving the highest pass@1 accuracy while also delivering the most substantial reduction in CoT length on code-related tasks.
+Importantly, SEER substantially mitigates truncation and infinite reasoning loops.
+As a result, reasoning loops are mitigated by up to 96.8%, leading to improved inference efficiency and output stability.
 In summary, our main contributions are as follows:
-  • We conduct an empirical study on HumanEval to examine how excessively long CoTs affect coding tasks.
-The results show that overlong reasoning often causes truncation, loops, and latency, ultimately reducing accuracy and stability, motivating the need for effective CoT length control.
-  • To address this, we propose SEER, a lightweight CoT compression framework that learns concise reasoning patterns from self-generated outputs, effectively reducing truncation and infinite loops without external modules.
-  • Extensive experiments across software engineering and mathematical reasoning tasks demonstrate that SEER achieves an average 42.1% reduction in CoT length while maintaining or even improving pass@1 accuracy.
-  • Ablation and cross-domain studies further confirm that adaptive CoT filtering is the key driver of efficiency and accuracy, with BoN sampling providing complementary benefits, and that SEER generalizes well across different task domains.
+  • We present a systematic empirical study to analyze the impact of long CoT reasoning in code generation.
+Our results indicate that excessive CoT is common and truncation is an important factor affecting model performance.
+Most truncations are associated with reasoning loops, which lower accuracy, raise latency, and make generation unstable.
+  • We propose SEER, a lightweight and self-enhancing CoT compression framework that makes CoT reasoning both effective and efficient for software engineering tasks.
+By learning concise reasoning patterns from self-generated outputs, SEER effectively suppresses reasoning loops and truncation without relying on external compression tools or manual annotations.
+  • We evaluate SEER across multiple software engineering tasks.
+Experimental results show that SEER achieves an average 41.6% reduction in CoT length while preserving or even improving pass@1 accuracy, largely due to improved robustness against truncation and overlong reasoning.
 Paper Organization.
@@
 2. Background
-2.1. Chain-of-Thought Efficiency
+2.1. The Efficiency of CoT
 LLMs have achieved remarkable success in complex reasoning tasks, largely due to advanced prompting techniques like CoT (Wei et al., 2022).
@@
 Consequently, a key research direction in CoT efficiency is to strike an optimal balance: retaining the reasoning benefits of CoT while minimizing its associated costs.
-Approaches in this area typically explore methods such as pruning redundant steps, learning to generate more concise reasoning paths (Kang et al., 2025; Xia et al., 2025), or selectively applying CoT based on task complexity (Wang et al., 2025a).
-2.2. Overthinking Phenomenon in CoT
+Approaches in this area typically explore methods such as pruning redundant steps, learning to generate more concise reasoning paths (Kang et al., 2025; Xia et al., 2025), or selectively applying CoT based on task complexity (Wang et al., 2025).
+2.2. Existing Issues in CoT
+Overthinking Phenomenon in CoT.
 Beyond simple inefficiency, the relationship between CoT length and accuracy is not a straightforward “longer is better.” Prior work shows that longer reasoning does not always translate into higher accuracy (Xie et al., 2025).
@@
 This decline is a direct result of overthinking, where excessive reasoning introduces noise, deviates from the correct reasoning path, and ultimately undermines accuracy (Chen et al., 2024a).
-Taken together, these findings highlight that the key challenge is to carefully control CoT length so as to reduce overthinking while preserving accuracy.
-2.3. Self-Generated Data for Self-Supervised Fine-Tuning
-Recent research has demonstrated that LLMs can bootstrap their own capabilities by generating candidate data, applying automated quality filtering, and then fine-tuning on the filtered subset.
-STaR illustrates this approach.
-The model generates CoT rationales and answers, discards incorrect outputs, and fine-tunes iteratively on the correct rationale–answer pairs, reaching reasoning performance comparable to models that are thirty times larger (Zelikman et al., 2022).
-Additionally, extensions such as Quiet-STaR generalize STaR by generating token-level rationales during continued pretraining, improving zero-shot reasoning on benchmarks like GSM8K and CommonsenseQA without explicit task fine-tuning (Zelikman et al., 2024).
-These methods collectively form a cohesive pipeline: self-generation, automatic filtering, and self-fine-tuning.
-This structure allows LLMs to independently improve their reasoning, domain knowledge, and generalization capabilities without the need for human-labeled data.
+Overall, these findings highlight that the key challenge is to carefully control CoT length to reduce overthinking while preserving accuracy.
+Looping Problem in CoT.
+Recent studies have revealed looping problems in CoT reasoning (Pipis et al., 2025), where LLMs repeatedly generate identical or semantically redundant reasoning segments instead of making progress toward a solution.
+This phenomenon also causes practical issues in real-world applications, as users have reported failures induced by looping behavior on various developer forums and agent frameworks (Significant Gravitas, 2023a; LangChain AI, 2024; Significant Gravitas, 2023b).
+Looping is particularly prominent under greedy or low-temperature decoding and has been observed across a wide range of modern reasoning models.
+Prior work suggests that this behavior stems from misestimated action probabilities during training: when advancing to the correct next reasoning step is difficult, models may favor repetitive continuations that are easier to predict, resulting in self-reinforcing loops.
+Such errors are further amplified by deterministic decoding, allowing repetition to persist across multiple steps.
+Although increasing sampling randomness can mitigate looping to some extent, it does not address the underlying learning deficiencies, indicating that looping is a fundamental inefficiency of CoT reasoning rather than a purely inference-time artifact.
 3. Empirical Study
-Before introducing our proposed framework, we conduct an empirical study to investigate how the length of CoT reasoning influences the performance, efficiency, and stability of LLMs in software engineering scenarios.
-Our analysis primarily targets coding tasks, a core application area of software engineering where precise and executable reasoning is essential.
+We first conduct an empirical study to investigate how the length of CoT reasoning influences the performance, efficiency, and stability of LLMs in software engineering scenarios.
+We choose a representative problem in software engineering, i.e., code generation tasks, as the subject of our empirical study.
 This study provides a quantitative assessment of the challenges introduced by excessively long CoT outputs, thereby motivating the design of our method.
@@
 3.1.1. Datasets
-Our empirical study is conducted on the HumanEval (et al., 2021), which is a widely used benchmark consisting of manually curated Python programming problems.
-Each problem contains a function signature, a docstring, and a set of test cases for evaluating functional correctness.
-Additionally, in the final part of our empirical study, we use a more complex dataset, Codeforces (Penedo et al., 2025).
-Codeforces comprises challenging algorithmic optimization problems collected from one of the most popular websites among competitive programmers.
+Our empirical study is conducted on two widely used benchmarks, i.e., HumanEval (et al., 2021) and MBPP-Sanitized (Austin et al., 2021).
+They consist of Python programming problems, each accompanied by a function signature, a natural-language docstring describing the task, and a set of unit tests used to evaluate functional correctness..
 3.1.2. LLMs
-We employ the DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025a) series of models to conduct our empirical study.
-This choice is motivated by the fact that it is one of the most popular reasoning models in recent times.
-Moreover, it is open-source, which facilitates deployment.
-We conduct experiments on four LLMs with sizes of 1.5B, 7B, 14B, and 32B.
+In our empirical study, we evaluate CoT behavior using a set of open-source reasoning-capable LLMs that are tested under the same experimental settings.
+The models cover different parameter scales and architectures, including the DeepSeek-R1-Distill-Qwen family (7B, 14B, and 32B), DeepSeek-R1-Distill-Llama-8B, as well as other widely used reasoning models such as QwQ-32B, Qwen3-8B, and gpt-oss-20b.
 3.1.3. Evaluation Metrics
-To assess the performance of CoT, we employ two metrics.
-[…]
-Accuracy.
-We use Pass@1 as the primary metric to evaluate correctness.
-This metric calculates the proportion of problems for which the model produces a correct solution on the first attempt, directly reflecting its practical effectiveness in solving tasks.
-[…]
+To assess the performance of CoT on software engineering tasks, we employ three metrics.
+Pass@1.
+Pass@1 calculates the proportion of problems for which the model produces a correct solution on the first attempt, directly reflecting its practical effectiveness in solving tasks.
+It is widely used on code generation tasks as it emphasizes single-attempt correctness under realistic deployment scenarios (et al., 2021).
 Average Token Length.
 We report the average number of tokens generated in model outputs.
-This metric reflects the verbosity of the reasoning process and helps assess whether excessively long outputs introduce redundancy or cause truncation.
-It also facilitates a direct comparison of output lengths between models with and without CoT reasoning.
+It reflects the verbosity of the reasoning process and helps assess whether excessively long outputs introduce redundancy or cause truncation.
+It also facilitates a direct comparison of output lengths among models with and without CoT reasoning.
+Truncation Number.
+We count the number of instances where generation is terminated due to exceeding the maximum token budget.
+This metric reflects the risk of uncontrolled or redundant reasoning, which may prevent the model from completing valid solutions and thus degrade real-world usability.
 3.1.4. Implementation Details
-[…]
+To ensure a fair and reliable empirical evaluation, we describe the implementation details of our experimental setup in this subsection.
+Disabling CoT.
+In our experiments, we evaluate models with explicit CoT disabled.
+Following the practice adopted by Qwen3-8B (Team, 2025), we suppress intermediate reasoning by appending an empty thinking segment, i.e., “<think> </think>”, to the input prompt.
+This operation prevents the model from emitting explicit reasoning traces while preserving its ability to produce final answers.
+Context Length and Prompt Control.
+To ensure fair comparison, the maximum context length is fixed to 16k tokens for all models across all experiments.
+To avoid prompt-induced variation in reasoning length, we exclude any instructions that explicitly encourage or constrain reasoning behavior (e.g., “Please reason step by step”).
+For example, for the HumanEval dataset, we use the following instruction:
+Please write a python program to address the following QUESTION. Your ANSWER should be in a code block format like this: ```python # Write your code here ```.
+Loop Detection.
+Following Pipis et al.
+(2025), we identify looping behavior using an $n$-gram repetition criterion.
+A response is classified as looping if any $n$-gram appears at least $k$ times within the same output.
+Unless otherwise specified, we set $n=30$ and $k=20$ for all reasoning models, consistent with the original study.
+This criterion captures severe degenerative repetition patterns commonly observed under greedy or low-temperature decoding and is empirically robust to moderate variations of $n$ and $k$.
 3.2. Empirical Findings
-[…]
-Truncation Leads to Performance Degradation.
-Table 1 demonstrates that excessively long CoT outputs introduce severe truncation issues, which directly limit performance improvements.
-For the smallest model (DeepSeek-R1-Distill-Qwen-1.5B), accuracy drops sharply, with the pass@1 score falling from 60.36% to 46.34%.
-This loss coincides with an extreme average CoT length of 7,645 tokens.
-The underlying cause is frequent truncation, which removes critical parts of the solution.
-In fact, 63 out of 164 problems are truncated on this model.
-As the model size increases, the truncation problem becomes less frequent, and moderate gains emerge.
-The 7B and 14B models show performance gains with CoT, but these improvements are limited by long outputs and truncation.
-The 32B model still experiences truncation in about 8.5% of cases, which keeps its pass@1 unchanged at 90.24% with or without CoT.
-Table 1. HumanEval Results of Reasoning Models (With vs. Without CoT)
-Models HumanEval (With CoT) HumanEval (Without CoT)
-Pass@1 (%) Avg Tokens Trunc Loop Pass@1 (%) Avg Tokens Trunc Loop
-DeepSeek-R1-Distill-Qwen-1.5B 46.34 7,645 63 62 60.36 513 0 0
-DeepSeek-R1-Distill-Qwen-7B 80.49 4,210 24 22 70.12 474 0 0
-DeepSeek-R1-Distill-Qwen-14B 89.63 3,687 16 14 84.76 444 0 0
-DeepSeek-R1-Distill-Qwen-32B 90.24 3,340 14 12 90.24 474 0 0
-[…]
-Finding 1:  Excessive CoT length induces harmful truncation, especially in small-scale LLMs.
-Small-scale LLMs are more easily affected by long CoTs because their context is quickly filled.
-For larger models, truncation is less frequent but still reduces the benefit, showing that longer reasoning does not always improve results once the context limit is reached.
-[…]
-Reasoning loop results in truncation.
-We manually examine truncation cases to better understand their impact on model performance.
-Our analysis shows that the model often falls into repetitive reasoning loops, repeatedly generating the same fragment.
-For example, it may continually output phrases such as “Wait, if …” until the maximum context length is reached.
-This behavior consumes a large number of tokens and severely reduces efficiency.
-As shown in Table 1, reasoning loops account for most truncation cases, with an average rate of 94.0%.
-When CoT generation is disabled, this problem disappears entirely.
-Larger models alleviate the issue to some extent, but reasoning loops are not fully eliminated.
-These observations suggest that while CoT can enhance reasoning, uncontrolled generation often leads to loops and truncation, which harm both efficiency and reliability.
-Therefore, controlling CoT length is crucial to preserve its benefits without incurring instability.
-[…]
-Finding 2:  Reasoning loops are the dominant reason for truncation.
-They waste tokens, reduce efficiency, and persist across model scales, disappearing only when CoT is disabled.
-This highlights looping as a critical issue in CoT reasoning.
-[…]
-Is longer always better?
-While prior work (Jin et al., 2024) suggests that increasing the length of CoT reasoning improves accuracy, our findings reveal a different pattern.
-We conduct experiments on one of the most challenging problems in HumanEval (findmyway, 2024), namely HumanEval/129.
-Since the 1.5B model achieves almost no correct solutions in our experiments (pass@1 is nearly zero), we exclude it from further analysis.
-For each of the other models, we generate 10,000 responses and examine the distribution of CoT lengths, as shown in Figure 1, with the maximum token length set to 16,384.
-Contrary to the common assumption that longer reasoning consistently leads to better performance (Jin et al., 2024), our analysis reveals a clear pattern across all three models.
-Failed outputs are frequently associated with longer CoT lengths, while successful outputs are more concentrated around shorter lengths.
-For instance, in the 7B model, the median CoT length of failed generations (9,489 tokens) surpasses that of successful generations (8,296 tokens), a pattern that persists in both the 14B and 32B models.
-This suggests that excessively long reasoning chains may introduce redundancy, amplify error propagation, or even trigger context truncation, ultimately degrading performance.
-These results highlight that effective reasoning does not simply require “more thinking”, but rather appropriate reasoning lengths that balance sufficiency and conciseness.
+How does CoT perform?
+Table 1 shows the performance of CoT in different models.
+The results reveal that excessive CoT length is a common phenomenon across most evaluated models.
+Except for gpt-oss-20b, all models generate long reasoning traces, with average CoT lengths ranging from 2K to 4K tokens.
+Importantly, non-negligible truncation rates are observed on both HumanEval and MBPP-Sanitized, reaching up to 17.1% in some settings.
+The results suggest that excessive CoT verbosity is a common issue in modern reasoning models.
+It appears consistently across diverse architectures and model sizes, frequently exceeding generation budgets.
+Table 1. Performance Comparison on HumanEval and MBPP-Sanitized (Max Generated Tokens = 16,384)
+Models HumanEval (164 Questions) MBPP-Sanitized (378 Questions)
+$pass@1(\%)$ CoT / Total Truncation # $pass@1(\%)$ CoT / Total Truncation #
+Qwen3-8B (w/o CoT) 84.76% - / 249 0 (0.0%) 82.01% - / 78 0 (0.0%)
+Qwen3-8B 89.02% 3,604 / 3,701 11 (6.7%) 84.13% 3,663 / 3,712 43 (11.4%)
+DeepSeek-Qwen-7B (w/o CoT) 65.24% - / 412 0 (0.0%) 50.79% - / 173 0 (0.0%)
+DeepSeek-Qwen-7B 78.04% 4,099 / 4,487 28 (17.1%) 73.54% 2,459 / 2,806 33 (8.7%)
+DeepSeek-Qwen-14B (w/o CoT) 82.32% - / 456 0 (0.0%) 74.87% - / 424 0 (0.0%)
+DeepSeek-Qwen-14B 87.20% 3,282 / 3,669 18 (11.0%) 86.51% 2,078 / 2,453 17 (4.5%)
+DeepSeek-Qwen-32B (w/o CoT) 87.20% - / 183 0 (0.0%) 88.62% - / 199 0 (0.0%)
+DeepSeek-Qwen-32B 88.41% 3,136 / 3,539 17 (10.4%) 88.10% 2,333 / 2,686 28 (7.4%)
+DeepSeek-Llama-8B 76.22% 3,640 / 4,021 19 (11.6%) 63.23% 2,568 / 2,910 35 (9.3%)
+QwQ-32B 93.90% 2,363 / 2,458 2 (1.2%) 93.65% 1,929 / 1,990 6 (1.6%)
+gpt-oss-20b 92.68% 481 / 790 1 (0.6%) 94.44% 470 / 693 3 (0.8%)
+    DeepSeek-Llama-8B, QwQ-32B and gpt-oss-20b do not support disabling CoT.
+To evaluate the impact of CoT, we compare the results with and without CoT on models that support disabling CoT reasoning.
+Enabling CoT improves pass@1 accuracy by an average of 7.4% across the evaluated models, indicating that CoT effectively enhances reasoning performance.
+Specifically, the performance gains are substantially larger for smaller models (e.g., +12.8% on HumanEval for DeepSeek-Qwen-7B), indicating that CoT plays a more critical role when model capacity is limited.
+However, these accuracy improvements come at a substantial efficiency cost.
+Enabling CoT increases the token cost by up to an order of magnitude in some cases (e.g., from approximately 400 tokens to over 4K tokens for DeepSeek-Qwen-7B), leading to significantly higher inference latency.
+Moreover, the resulting long CoT frequently triggers truncation, which partially offsets the accuracy benefits of CoT.
+In several settings, truncated CoT leads to degraded performance compared to non-CoT generations, particularly when the final answer is cut off.
+For example, on MBPP-Sanitized, DeepSeek-Qwen-32B shows a small decrease in pass@1 (from 88.62% to 88.10%), alongside a truncation rate of 7.4%.
+Overall, these results highlight a fundamental trade-off in the use of CoT.
+While CoT improves reasoning accuracy, its effectiveness is constrained by excessive verbosity, increased latency, and truncation-induced failures.
+This observation motivates a closer look at why CoT becomes excessively long and how much of the generated reasoning is truly needed for solving the problem, which we explore in the next sections.
+    Finding 1:  CoT improves accuracy but induces excessive verbosity and truncation.
+Across various models, enabling CoT improves pass@1 accuracy by an average of 7.4%.
+However, it can increase token cost about 10 times and frequently cause truncation, with a maximum rate of 17.1%.
+How Common is Looping among Truncations?
+To quantify how often truncation is associated with looping behavior, we analyze all truncated outputs using the loop detection criterion described in Section 3.1.4.
+Across all 261 truncation cases observed across different models and datasets, 236 cases (90.4%) are identified as looping by the $n$-gram repetition criterion.
+Since our loop detector is intentionally strict (n=30n{=}30, k=20k{=}20), borderline cases with weaker or shorter repetition may be missed.
+We manually examine the remaining truncation cases not flagged by the detector and find that some of them still exhibit loop-like repetition patterns, although they do not satisfy the strict $n$-gram threshold.
+Therefore, the 91.8% should be viewed as a conservative lower bound, and the true fraction of truncations affected by looping is likely higher.
+    Finding 2:  Looping is highly common in truncated outputs.
+Using a strict $n$-gram repetition detector, 91.8% of truncations are identified as looping, and manual inspection suggests this is a conservative estimate.
+How Much Reasoning is Effective?
+To quantitatively assess the proportion of useful reasoning within CoT, we conduct a LLM-based CoT compression.
+Following prior work that employs strong LLMs as judges or analyzers to evaluate model-generated text (Zheng et al., 2023), we use GPT-4o (OpenAI, 2024) to extract the effective reasoning steps in CoT outputs.
+Specifically, this process removes the surface-level verbosity while preserving the original logic.
+Essential reasoning steps and planning instructions are retained.
+The prompt used for this section is provided in the replication package.
+To reduce randomness in LLM-based compression, we perform the compression three times for each CoT and report the averaged results.
+After compression, we measure 1) the token length of the compressed CoT and 2) the effective reasoning ratio, defined as the fraction of tokens retained after compression relative to the original CoT length.
+The results are reported in Table 2.
+Truncated instances are omitted because they have not finished the reasoning process.
+Table 2. Effective Reasoning Length and Retention Ratio Estimated via GPT-4o on HumanEval and MBPP
+Models HumanEval (164 Questions) MBPP-Sanitized (378 Questions)
+Avg. CoT (Before) Avg. CoT (After) Retained (%) Avg. CoT (Before) Avg. CoT (After) Retained (%)
+DeepSeek-Llama-8B 3,594 212 5.91 2,538 164 6.50
+DeepSeek-Qwen-7B 3,677 210 5.71 2,334 177 7.62
+DeepSeek-Qwen-14B 3,076 211 6.88 1,974 181 9.17
+DeepSeek-Qwen-32B 2,869 212 7.42 2,218 179 8.07
+QwQ-32B 2,311 216 9.36 1,866 187 10.05
+Qwen3-8B 3,456 228 6.62 3,527 155 4.41
+gpt-oss-20b 481 149 31.11 470 128 27.27
+Table 2 shows significant redundancy in generated CoTs.
+For most models (except gpt-oss-20b), the compressed lengths consistently fall between 150 and 230 tokens across benchmarks.
+Consequently, the effective reasoning ratio falls below 10%, suggesting that most generated tokens are unnecessary for the core reasoning.
+Additionally, gpt-oss-20b shows more efficient reasoning, but approximately 70% of its reasoning is still ineffective.
+A closer comparison between Qwen3-8B and QwQ-32B on MBPP-Sanitized further illustrates this phenomenon.
+Although Qwen3-8B generates nearly twice as many CoT tokens as QwQ-32B before compression, its effective reasoning ratio after compression is even lower.
+This indicates that increased CoT verbosity does not correspond to more effective reasoning and may instead amplify redundant or low-value content.
+Overall, these results indicate that longer CoT is not necessary, and motivate the need for CoT compression that focuses on preserving essential reasoning steps.
+    Finding 3:  Only a small fraction of CoT contributes to essential reasoning.
+Compression-based analysis with GPT-4o shows that, despite large variation in raw CoT length, less than 10% of CoT tokens are retained across most models.
+The results indicate substantial redundancy in generated reasoning.
+Is Longer Reasoning Always Better?
+To evaluate whether longer CoT reasoning leads to better outcomes, we generate multiple responses for the same problem and analyze the CoT lengths of passed and failed solutions.
+We focus on a relatively difficult problem, where the quality of CoT is more likely to influence the final result.
+Since HumanEval does not provide an official difficulty classification, we select the problem with the longest prompt in the HumanEval dataset, i.e., HumanEval/129.
+We choose the DeepSeek-Qwen-Series Model for our evaluation, focusing on how the length of CoT impacts its performance across various model sizes.
+The prompt follows the implementation details of this section (Section 3.1.4), without any explicit instructions to control the length of the CoT.
+For each of the models, we generate 10,000 responses and examine the distribution of CoT lengths, as shown in Figure 2, with the maximum token length set to 16,384.
 Figure 2.
@@
 An excessively long reasoning chain can instead lead to a decline in accuracy.
+While prior work (Jin et al., 2024) suggests that increasing the length of CoT reasoning improves accuracy, our findings reveal a different pattern across all three models.
+Failed outputs are frequently associated with longer CoT lengths, while successful outputs are more concentrated around shorter lengths.
+For instance, in the 7B model, the median CoT length of failed generations (9,489 tokens) surpasses that of successful generations (8,296 tokens), with a similar pattern observed in both 14B and 32B models.
+A closer examination of failure cases shows that overly long reasoning often introduces redundancy, propagates errors, and may lead to truncation, which ultimately degrades performance.
+These results highlight that effective reasoning does not simply require “more thinking”, but rather appropriate reasoning lengths that balance sufficiency and conciseness.
+    Finding 4:  Longer CoT does not guarantee higher accuracy and is often associated with failures.
+On HumanEval/129, failed cases consistently show longer CoT than successful ones across all models (e.g., 9,489 vs.
+8,296 tokens in 7B), indicating that excessive reasoning is correlated with failures.
+Can Prompt Design Effectively Reduce CoT Length?
+Prior work has shown that the verbosity of CoT reasoning is sensitive to instruction-level guidance (Wei et al., 2022; Wang et al., 2023; Agrawal et al., 2024).
+Therefore, to verify whether the prompt design can effectively shorten the length of the CoT, we include a lightweight prompting baseline to encourage concise reasoning.
+Specifically, we augment the original task instruction with a general conciseness guideline, asking the model to provide its reasoning as concisely as possible while avoiding redundant or unnecessary steps.
+Concretely, we add the following system prompt:
+Please provide your reasoning as concisely as possible, avoiding redundant or unnecessary steps.
+Table 3. Effect of Conciseness-Guided Prompting on CoT Length and Model Accuracy
+Models HumanEval (164 Questions) MBPP-Sanitized (378 Questions)
 […]
-Finding 3:  Longer CoT does not guarantee higher accuracy.
-Across all three models, failed generations consistently exhibit longer CoT lengths than successful ones (Figure 1).
-This indicates that excessively long reasoning chains often introduce redundancy and error propagation rather than improving performance, highlighting the need for appropriately controlled CoT lengths.
-[…]
-Long CoT introduces application latency.
-Excessively long CoT reasoning introduces significant latency overhead, which ultimately degrades user experience.
-To quantify this impact, we evaluate the inference performance of Deepseek-R1-Distill-Qwen models of varying sizes using the transformers library on a single NVIDIA A800 GPU.
-The measured generation speeds range from 46 tokens/s for the smallest (1.5B) model to 18 tokens/s for the largest (32B) model.
-Using these measurements, we estimate the average end-to-end latency on two benchmarks: HumanEval, representing relatively simple coding tasks, and Codeforces, which contains more complex algorithmic challenges.
-The results are presented in Figure 2.
-Figure 2.
-Inference latency of different model sizes on HumanEval and Codeforces datasets with and without CoT reasoning.
-While CoT improves reasoning capability, it substantially increases application latency, particularly for larger models, posing practical constraints on real-world deployment.
-Figure 2 shows that incorporating CoT substantially increases response time, particularly for larger models on complex tasks.
-On Codeforces, the 32B model with CoT requires an average of 776 seconds to complete the reasoning for each problem, compared to 157 seconds without CoT.
-Even the smaller 1.5B model takes more than five minutes per problem when using CoT.
-These delays make real-time use impractical.
-Moreover, all experiments were conducted under a 16k-token context limit, whereas Codeforces samples already exceed 14k tokens on average, implying that real-world latency would be even higher.
-Thus, while CoT improves reasoning, its verbosity imposes severe constraints on latency-sensitive applications, highlighting the need for effective CoT length control.
-[…]
-Finding 4:  Long CoT substantially slows down inference.
-For example, the 32B model on Codeforces requires nearly five times longer with CoT than without.
-This latency overhead makes CoT impractical for real-time or latency-sensitive applications, underscoring the need for effective length control.
+DeepSeek-Llama-8B 3,640 3,063 (-15.9%) +2.44% 2,568 2,564 (-0.2%) -1.85%
+DeepSeek-Qwen-7B 3,839 3,464 (-9.8%) +0.61% 2,459 1,946 (-20.9%) -5.55%
+DeepSeek-Qwen-14B 3,282 2,890 (-11.9%) +1.82% 2,078 1,952 (-6.1%) -1.06%
+DeepSeek-Qwen-32B 3,136 2,610 (-16.8%) +1.83% 2,333 1,890 (-19.0%) +1.58%
+QwQ-32B 2,363 2,236 (-5.4%) 0.00% 1,929 1,779 (-7.8%) +0.53%
+Qwen3-8B 3,604 2,490 (-30.9%) +3.05% 3,663 2,545 (-30.5%) +3.70%
+gpt-oss-20b 481 238 (-50.5%) +4.88% 470 241 (-48.7%) -2.64%
+As shown in Table 3, introducing this prompt instruction reduces the average CoT length for most evaluated models.
+This confirms that CoT verbosity is sensitive to instruction-level guidance.
+However, the reduction rate vary substantially across models and datasets, indicating that prompt-based control lacks reliability and offers limited effectiveness in practice.
+In particular, while moderate reductions are observed for several DeepSeek and Qwen models, the overall compression ratios are relatively small and highly model-dependent.
+In contrast, gpt-oss-20b exhibits a much more pronounced reduction in reasoning length (over 50% on HumanEval).
+One possible explanation is that gpt-oss-20b natively supports explicit control of reasoning length, whereas most other open-source models lack such mechanisms, making them less responsive to high-level conciseness instructions.
+It should be noted that shortening CoT via prompting does not uniformly degrade performance.
+In several cases, accuracy remains stable or even improves after applying the concise-CoT prompt (e.g., Qwen3-8B and DeepSeek-Qwen-32B on HumanEval).
+This observation reinforces our Finding 3.2 that excessively long CoT does not necessarily lead to better reasoning.
+On the contrary, overly verbose reasoning may instead introduce redundant or noisy content that hinders effective problem-solving.
+Overall, prompt design can influence CoT length but lacks robustness and consistency across models.
+The fact that shorter CoT can maintain or even improve accuracy shows that reducing unnecessary reasoning is both feasible and useful.
+    Finding 5:  Prompt-based conciseness control is limited and model-dependent.
+Conciseness prompts reduce CoT length for most models, but the gains are generally small and inconsistent.
+Only gpt-oss-20b shows a substantial reduction (over 50% on HumanEval), while accuracy is largely preserved.
 4. Methodology
@@
 In this step, the base language model is used to process each question in the training set, generating a complete solution that includes an explicit, step-by-step reasoning path.
-A distinctive advantage of our approach is its low sensitivity to prompt engineering.
-While a prompt is required to elicit CoT reasoning, our subsequent adaptive filtering stage ensures that the method is not critically dependent on the choice of prompt format.
-This property allows us to adopt a pragmatic strategy: whenever an official, dataset-specific prompt template is available, we use it directly; otherwise, we revert to a robust general-purpose template.
-For example, for mathematical reasoning tasks, we utilize the system prompt from the study by Wei et al.
-(Wei et al., 2022):
+This step does not rely on prompt design.
+Therefore, we intentionally avoid using complex prompts.
+For example, for code generation tasks, we adopt the same prompt as in Section 3.1.4.
+To avoid early truncation that may prevent the model from completing its reasoning and based on the reasoning length observed in our empirical study, we adopt a moderate token budget during generation, i.e., 16k.
+Moreover, this setting can help collect more complete and diverse reasoning traces, thereby reducing information loss and mitigating overfitting caused by insufficient CoT data.
+4.1.1. Best-of-N Sampling for Data Refinement
+Since reasoning loops are a common phenomenon in CoT generation for reasoning models, we adopt a BoN sampling strategy to mitigate looping behaviors.
+In addition, this strategy enables us to collect more diverse reasoning traces for constructing a high-quality training set, and can be used to expand the dataset when the available training data are limited.
+Specifically, for each input question, the model generates N candidate responses, each accompanied by a CoT.
+These candidates are then ranked and filtered according to a set of criteria: ❶ Correctness of the Final Answer: only candidates producing the correct final answer are considered.
+All incorrect outputs are discarded.
+❷ Presence of Valid CoT: from the remaining candidates, we retain only those containing a non-empty and non-looping reasoning path.
+This guarantees that each example is explainable and preserves its interpretability.
+❸ Conciseness of Reasoning: if multiple candidates satisfy the first two criteria, we select the one with the shortest CoT length.
+This prioritizes efficiency and prevents the inclusion of unnecessarily verbose reasoning.
+4.2. Adaptive CoT Filtering
+After BoN sampling, CoT lengths remain highly variable and often contain long-tail verbosity. To obtain concise yet correct training signals, we apply a two-stage filter.
+Stage 1: Output validity.
+We first remove samples whose final outputs are invalid (e.g., code that cannot be parsed or fails the provided unit tests in code-generation datasets).
+This step filters out malformed or incomplete outputs before applying any length-based filtering.
+Stage 2: Robust length filter.
+On the remaining samples, we filter overly long CoTs using a robust, distribution-aware threshold.
+This is motivated by the empirical observation that CoT lengths are often heavy-tailed: a small fraction of samples can be extremely long due to overthinking or looping, and these outliers can dominate both training cost and the learned behavior.
 […]
-To prevent incomplete reasoning caused by premature output truncation, we allocate a sufficiently high maximum token budget during generation.
-This ensures that the model’s natural CoT length distribution is faithfully captured.
-Accurately profiling this distribution is critical for calibrating the adaptive filter in later stages, as it allows us to set dataset-specific compression thresholds that minimize information loss.
-4.2. Best-of-N Sampling for Data Refinement
-To prevent the reasoning loops noted in Section 3, which can harm data quality, we adopt a BoN sampling strategy to build a high-quality training set.
-This strategy also helps obtain cleaner datasets with more concise CoT lengths.
-For each input question, the model generates $\mathcal{N}$ candidate responses, each augmented with a CoT.
-These candidates are then ranked and filtered according to a strict hierarchical set of criteria:
-  1. (1)
-    Correctness of the Final Answer: Only candidates producing the correct final answer are considered. All incorrect outputs are discarded.
-  2. (2)
-    Presence of Valid CoT: From the remaining candidates, we retain only those containing a non-empty reasoning path.
-This guarantees that each example is explainable and preserves its interpretability.
-  3. (3)
-    Conciseness of Reasoning: If multiple candidates satisfy the first two criteria, we select the one with the shortest CoT length.
-This prioritizes efficiency and prevents the inclusion of unnecessarily verbose reasoning.
-This multi-stage filtering process ensures that the resulting dataset consists exclusively of reasoning traces that are (1) correct in outcome, (2) logically explicit, and (3) concise.
-By integrating BoN sampling into our pipeline, we are able to suppress reasoning loops, improve data quality, and provide the model with more effective fine-tuning signals.
-4.3. Adaptive CoT Filtering
-Even after BoN sampling, CoT responses can vary widely in length depending on the task, domain, and problem complexity.
-Excessively long CoTs waste computational resources, while overly aggressive compression risks discarding essential reasoning steps.
-To balance these factors, we introduce an adaptive CoT filtering mechanism that sets a dataset-specific maximum CoT length, denoted $\lambda_{c}$.
-A fixed, manually chosen $\lambda_{c}$ is often suboptimal because optimal CoT lengths differ substantially across domains (e.g., mathematical proofs vs.
-code debugging).
-If $\lambda_{c}$ is too low, important intermediate reasoning may be lost, potentially reducing accuracy and even causing the model to overfit to overly compressed reasoning patterns.
-Conversely, a value set too high will dilute the benefits of compression, leading to minimal gains in efficiency.
-Our adaptive filtering addresses this by first generating CoT responses for the entire training set during the pre-inference stage and then computing two statistical measures of their length distribution: the mean length $\bar{\lambda}$ and the median length $\tilde{\lambda}$.
-We define $\lambda_{ada}$ as:
-$\lambda_{ada}=\frac{\bar{\lambda}+\tilde{\lambda}}{2}$
-This formula anchors the threshold to the observed characteristics of each dataset, balancing robustness (median) with representativeness (mean).
-While $\lambda_{c}$ provides a principled baseline, our design remains flexible.
-Users can manually adjust it to match specific efficiency or accuracy requirements.
-This tunable control allows practitioners to trade off between computational savings and reasoning completeness, enabling SEER to adapt seamlessly across diverse reasoning scenarios.
-4.4. Overall Procedure
-Input: Training set $\mathcal{D}=\{(x_{i},y_{i}^{\star})\}$, base model $\mathcal{M}$, candidates per input $\mathcal{N}$
-Output: The model $\mathcal{M}_{\text{adapt}}$ fine-tuned on the filtered dataset $\widetilde{\mathcal{D}}$
-Stage 1: Pre-inference generation;
-foreach $(x_{i},y_{i}^{\star})\in\mathcal{D}$ do
-  Generate $\mathcal{N}$ CoT-augmented candidates with $\mathcal{M}$ under a fixed prompt;
-  $r_{ij}\leftarrow\text{final response generated for }x_{i}$;
-  $\text{CoT}_{ij}\leftarrow\text{reasoning trace (Chain-of-Thought) leading to }r_{ij}$;
-  $\mathcal{C}_{i}\leftarrow\{\,(r_{i1},\text{CoT}_{i1}),\ldots,(r_{i\mathcal{N}},\text{CoT}_{i\mathcal{N}})\,\}$;
-Stage 2: Best-of-N (BoN) sampling;
-foreach $(x_{i},y_{i}^{\star}),\mathcal{C}_{i}$ do
-  $\mathcal{S}_{i}\leftarrow\{(r,\text{CoT})\in\mathcal{C}_{i}\mid\textsc{Correct}(r,y_{i}^{\star})\ \wedge\ |\text{CoT}|>0\}$ if $\mathcal{S}_{i}\neq\emptyset$ then
-     $(r_{i}^{\star},\text{CoT}_{i}^{\star})\leftarrow\arg\min_{(r,\text{CoT})\in\mathcal{S}_{i}}|\text{CoT}|$ Add $(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})$ to $\mathcal{D}^{\prime}$;
-Stage 3: Adaptive CoT filtering;
-Compute length set $L\leftarrow\{\,|\text{CoT}_{i}^{\star}|\mid(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})\in\mathcal{D}^{\prime}\,\}$;
-$\bar{\lambda}\leftarrow\text{mean}(L),\ \tilde{\lambda}\leftarrow\text{median}(L)$;  $\lambda_{c}\leftarrow(\bar{\lambda}+\tilde{\lambda})/2$;
-$\widetilde{\mathcal{D}}\leftarrow\{(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})\in\mathcal{D}^{\prime}\mid|\text{CoT}_{i}^{\star}|\leq\lambda_{c}\}$;
-Stage 4: Fine-tuning;
-Fine-tune model on $\widetilde{\mathcal{D}}$ with SFT to obtain $\mathcal{M}_{\text{adapt}}$;
-return $\mathcal{M}_{\text{adapt}}$;
-Algorithm 1 Self-Enhancing Training Process with BoN & Adaptive CoT Filtering
-Algorithm 1 summarizes the end-to-end process, from pre-inference generation to BoN-based selection, adaptive filtering, and final fine-tuning on the curated dataset.
-Through the combination of high-coverage pre-inference generation, strict BoN filtering, and adaptive CoT length control, SEER produces training data that is both high-quality and efficiency-oriented.
-This not only improves the reasoning efficiency of the fine-tuned model but also mitigates failure modes such as infinite reasoning loops and truncation.
 5. Experimental Setup
 5.1. Datasets
-To thoroughly evaluate the performance and efficacy of our method, we conduct experiments across three distinct software engineering tasks and one mathematical reasoning task.
-These tasks were selected to cover a broad spectrum of problem types, input–output modalities, and reasoning requirements.
-By incorporating both code-related and mathematical reasoning benchmarks, we aim to investigate not only the in-domain effectiveness of our method but also its adaptability to different reasoning paradigms and task formats.
-  • MathQA-Python (Austin et al., 2021) is a dataset designed for generating Python programs that solve mathematically formulated problems.
-Each problem is presented in natural language and requires the model to produce an executable Python solution.
-The challenges in this dataset include parsing problem statements, performing intermediate symbolic reasoning, and implementing the reasoning steps as correct Python code.
-This dataset serves as a baseline for evaluating code generation in mathematically oriented problem-solving contexts.
-  • CodeXGLUE-Defect-Detection (Zhou et al., 2019) is a binary classification dataset that assesses a model’s ability to identify whether a given piece of source code contains security vulnerabilities or functional defects.
-The data originates from real-world code bases, and the detection task often requires understanding complex control flows, data dependencies, and security implications.
-This dataset evaluates the model’s capability in static code analysis and secure software engineering scenarios, where concise yet accurate reasoning is critical to avoid excessive inference overhead.
-  • Code-Search-WebQuery (Lu et al., 2021; Husain et al., 2019) is used to test semantic code retrieval capabilities.
-In this task, the model is provided with a natural language query and a code snippet, and must determine whether the code correctly satisfies the described requirement.
-This problem is framed as binary classification, and it challenges the model to bridge the semantic gap between natural language descriptions and programming logic.
-The task reflects realistic software development workflows, such as searching for relevant functions in large code bases.
-  • GSM-8K (Cobbe et al., 2021) is a large-scale benchmark for multi-step mathematical reasoning.
-Each question requires systematic decomposition into intermediate steps before arriving at the final answer.
-Unlike MathQA-Python, the output here is not executable code but a final numeric answer, making it an ideal testbed for evaluating whether our method can compress reasoning chains without degrading accuracy in purely symbolic reasoning scenarios.
-It also provides an out-of-domain evaluation for models fine-tuned on code-related tasks.
-Additionally, in RQ3 we include the HumanEval dataset introduced in Section 3 to evaluate the generalizability of our method.
-By selecting this combination of datasets, our evaluation covers tasks ranging from code synthesis, defect detection, and semantic code retrieval to mathematical reasoning.
+To evaluate the performance and effectiveness of our method, we conduct experiments on a set of representative software engineering datasets.
+The experiments covers diverse task formats and reasoning requirements.
+5.1.1. Training and evaluation benchmarks.
+We use MathQA-Python (Austin et al., 2021), CodeXGLUE-Defect-Detection (Zhou et al., 2019), and Code-Search-WebQuery (Lu et al., 2021; Husain et al., 2019).
+All of these datasets provide predefined training and test splits.
+MathQA-Python focuses on generating executable Python programs to solve mathematically formulated problems, requiring both symbolic reasoning and correct code implementation.
+CodeXGLUE-Defect-Detection is a binary classification task that assesses whether a given code snippet contains functional defects or security vulnerabilities, emphasizing program understanding and static reasoning.
+Code-Search-WebQuery evaluates semantic code retrieval by determining whether a code snippet satisfies a natural language query, challenging the model to align program semantics with textual descriptions.
+These three datasets are used to train our method and all baseline models under identical settings, with evaluation conducted on their respective test sets.
+5.1.2. Evaluate-only benchmarks.
+In addition, we evaluate our method on HumanEval and MBPP-Sanitized, which are used exclusively for evaluation.
+These datasets are not involved in training, and serve to assess the generalization capability of our method on unseen code generation tasks.
+Overall, this benchmark suite enables a systematic evaluation of both in-domain performance and cross-task generalization across software engineering scenarios.
 5.2. Implementation Details
 All experiments are conducted on a server equipped with four NVIDIA RTX A800 GPUs (80 GB each) and 128 Intel Xeon CPUs, running Ubuntu 22.04 LTS.
-We fine-tune DeepSeek-R1-Distill-Qwen-7B model (DeepSeek-AI, 2025b) under a full-parameter supervised fine-tuning (SFT) setting, using the LLaMAFactory framework (Zheng et al., 2024).
-Training is performed for three epochs with an effective batch size of eight, gradient checkpointing enabled, and the AdamW optimizer.
+We fine-tune DeepSeek-R1-Distill-Qwen-7B model (DeepSeek-AI, 2025) under a full-parameter supervised fine-tuning (SFT) setting, using the LLaMAFactory framework (Zheng et al., 2024).
+Training is performed for three epochs with an effective batch size of eight, gradient checkpoint enabled, and the AdamW optimizer is used.
 The learning rate is fixed at $1\times 10^{-5}$ with a cosine decay schedule and 10% warm-up, and gradient clipping (norm 1.0) is applied for stability.
-During fine-tuning, inputs are tokenized with a maximum sequence length of 8,192 tokens, truncating from the input side if necessary to preserve the output reasoning chain.
+During fine-tuning, inputs are tokenized with a maximum sequence length of 16,384 tokens.
 Although we adopt full-parameter SFT by default, the SEER framework is also compatible with parameter-efficient fine-tuning methods such as LoRA and QLoRA, which we discuss in the Discussion section.
@@
 In our work, we consider the following baselines.
+  • Short CoT: Employ prompting techniques to elicit a more concise CoT from the model, using the prompt design described in Section 3.
   • Self-Training: Directly using self-inference data to train the model.
   • TokenSkip (Xia et al., 2025): Fine-tuning LLMs by selectively skipping less important tokens while retaining key tokens, enabling the generation of compressed CoTs with adjustable ratios.
-  • Naive BoN (Munkhbat et al., 2025b): Use only the Best-of-N method to process the pre-inference data and then train the model.
-In order to ensure the fairness of the experiment, we have implemented several modifications to the TokenSkip source code.
-These changes include: 1) the incorporation of the “deepseek” prompt template, as this template is employed by the chosen model; and 2) the use of the vllm library in place of the transformers library to achieve more efficient inference and to mitigate any potential precision loss that could lead to unfair comparisons.
+  • Naive BoN (Munkhbat et al., 2025): Use only the Best-of-N method to process the pre-inference data and then train the model.
 5.4. Evaluation Metrics
 We measure the following metrics:
-[…]
-Accuracy: We use the same pass@1, which is mentioned in Section 3 as the primary metric to evaluate correctness.
+Pass@1: We report pass@1 as the primary metric to evaluate correctness, i.e., whether the model produces a correct output on the first attempt.
 For programming tasks (e.g., MathQA-Python), a prediction is considered correct if the generated code passes all test cases.
-For mathematical reasoning tasks (e.g., GSM-8K), we adopt exact matching criterion, where the model’s final answer must exactly match the ground truth.
-[…]
+For binary classification tasks (e.g., CodeXGLUE-Defect-Detection), a prediction is considered correct if the predicted label matches the ground-truth label, and the resulting accuracy is equivalent to pass@1 under single-attempt evaluation.
 Compression Rate: We define the Compression Rate $\rho$ to quantify the conciseness of the generated CoT.
 Specifically, $\rho$ measures the relative reduction in the number of tokens compared to a reference CoT.
-It is computed as: $\rho=(\lambda-{\lambda^{\prime}})/{\lambda}$
+It is computed as $\rho=1-(\lambda^{\prime}/\lambda)$.
 In this formula, $\lambda$ denotes the token count of the reference (original) CoT, and $\lambda^{\prime}$ represents the token count of the CoT generated by the evaluated model or baseline.
 A higher value of $\rho$ indicates a greater degree of compression, corresponding to a more concise reasoning path.
-This metric is particularly important for evaluating methods such as TokenSkip and Naive BoN, which are explicitly designed to produce shorter CoTs without compromising correctness.
 6. Experiment
@@
   • RQ1: How does the CoT compression performance of SEER compare with baselines? We evaluate SEER against baseline methods on multiple datasets.
-  • RQ2: What is the contribution of each component of SEER?
-We conduct ablation studies by removing or varying parts of SEER, such as the CoT length filter $\lambda_{c}$ and the BoN parameter $\mathcal{N}$, to examine their impact.
-  • RQ3: How well does SEER generalize across domains? We examine whether SEER remains effective when the fine-tuned data and test data come from different domains.
-  • RQ4: Can SEER mitigate reasoning loops? We test whether SEER helps LLMs generate valid answers without falling into infinite reasoning loops.
+  • RQ2: How well does SEER generalize across domains? We examine whether SEER remains effective when the fine-tuned data and test data come from different domains.
+  • RQ3: What is the contribution of each component of SEER?
+We conduct ablation studies by removing or varying parts of SEER, such as the length-filter strictness $\alpha$ (which determines $\lambda_{c}$) and the BoN parameter $\mathcal{N}$, to examine their impact.
+  • RQ4: How can SEER mitigate reasoning loops? We test whether SEER helps LLMs generate valid answers without encounter looping problem.
 6.2. RQ1: The Performance of SEER
-In this subsection, we assess the performance of SEER in comparison with baseline methods.
-For the TokenSkip experiment, we set $\gamma=0.5$, meaning that we append the sequence “[EOS]0.5[EOS][EOS]0.5[EOS]” to the end of the input prompt.
-Similarly, for Naive BoN, we set $N=3$, which is consistent with the parameters used in evaluating our method.
-Table 2 reports task pass@1 accuracy ($\mathcal{A}$), average CoT length ($\lambda_{\text{avg}}$), and compression rate ($\rho$) under the 8K token constraint.
-Methods MathQA-Python Defect-Detection Code-Search GSM-8K $\rho_{avg}(\%)$
-$\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$
-Base 63.7 1,456 - 44.7 1,836 - 72.4 472 - 90.1 1,309 - -
-Self-Trained 74.6 1,177 19.2 49.3 1,055 42.5 72.5 468 0.9 89.2 792 39.5 25.5
-Naive BoN (N=3) 73.6 1,191 18.2 49.8 1,042 43.3 74.9 478 -1.3 89.8 845 35.5 23.9
-TokenSkip ($\gamma$=0.5) 1.6 1,196 17.9 36.2 2,421 -31.9 66.5 465 1.5 37.9 1,238 5.4 -1.8
-TokenSkip ($\gamma$=0.8) 0.9 1,109 23.8 47.2 1,004 45.3 76.4 341 27.8 87.3 1,197 8.6 26.4
-SEER 74.9 877 39.8 50.5 785 57.2 77.3 362 23.3 90.5 677 48.3 42.1
+In this subsection, we evaluate the effectiveness of SEER under the 16K token constraint by comparing it with baselines.
+For baselines with tunable parameters, we adopt commonly used and representative settings.
+Specifically, TokenSkip is evaluated with $\gamma=0.5$ (i.e., compression ratio) to examine its maximum compression behavior, and with $\gamma=0.8$ to reflect a more conservative and widely used configuration.
+For Naive BoN, we sample $N=3$ candidates, which matches the sampling budget used in SEER.
+For BoN section in SEER, we set the same parameter, i.e., $N=3$.
+The results are summarized in Table 4.
+Methods MathQA-Python Defect-Detection Code-Search $\rho_{avg}(\%)$
+$pass@1(\%)$ $\lambda_{avg}$ $\rho(\%)$ $pass@1(\%)$ $\lambda_{avg}$ $\rho(\%)$ $pass@1(\%)$ $\lambda_{avg}$ $\rho(\%)$
+Base 63.7 1,456 - 44.7 1,836 - 72.4 472 - -
+TokenSkip ($\gamma$=0.5) 1.6 1,196 17.9 36.2 2,421 -31.9 66.5 465 1.5 -4.2
+TokenSkip ($\gamma$=0.8) 0.9 1,109 23.8 47.2 1,004 45.3 76.4 362 23.3 30.8
+Self-Trained 74.6 1,177 19.2 49.3 1,055 42.5 72.5 468 0.9 20.9
+Naive BoN (N=3) 73.6 1,191 18.2 49.8 1,042 43.3 74.9 478 -1.3 20.1
+Short CoT 55.6 1,236 15.1 46.6 1,198 34.8 69.7 398 15.7 21.9
+SEER 74.9 877 39.8 50.5 785 57.2 77.3 341 27.8 41.6
     The Base model refers to “DeepSeek-R1-Distill-Qwen-7B”. The same applies to the other tables;
-    $\mathcal{A}(\%)$: Pass@1; $\lambda_{avg}$: Average output length (tokens);
-    $\rho(\%)$: Compression rate compared to the Base model; $\rho_{avg}(\%)$: Average compression rate across all datasets.
-Table 2. The CoT Compression Performance Over Different Datasets (8K Token Limit)
-Across all three software engineering tasks, SEER shows strong performance in both accuracy and CoT compression.
-On MathQA-Python, which requires code generation, it achieves 74.9% pass@1 with a 39.8% reduction in reasoning length.
-On Defect-Detection, it obtains the best pass@1 (50.5%) and the highest compression rate (57.2%).
-On Code-Search, it maintains a high pass@1 of 77.3% while compressing CoTs by 22.9%.
-Beyond software engineering, SEER also performs well on GSM-8K, a math reasoning dataset, reaching 90.5% pass@1 with a 48.3% reduction in CoT length.
-The accuracy gains are largely due to the mitigation of truncation and excessive reasoning, which often undermine model reliability.
-These results demonstrate that SEER consistently balances brevity and accuracy across diverse tasks.
-[…]
-Compared to TokenSkip, both Self-Trained and Naive BoN achieve more stable outcomes but still fall short of SEER.
-On Defect-Detection, their pass@1 remains below 50%, with compression rates under 30%.
-While they provide modest improvements over the base model, they cannot match SEER’s combination of high accuracy and substantial compression.
-This suggests that generic fine-tuning or simple BoN sampling is insufficient for the challenging trade-off between performance and efficiency.
-[…]
-Answer to RQ1: Compared with all baselines, SEER consistently achieves high accuracy and efficient CoT compression on both software engineering tasks and the math reasoning task.
-On average, SEER reduces CoT length by 42.1%, representing a 16.9% improvement over the best existing baseline, while preserving or even improving accuracy across diverse tasks.
-6.3. RQ2: The Impact of Each Components
-To investigate the contribution of each component, we conduct ablation studies on the MathQA-Python dataset by varying the BoN sampling size $\mathcal{N}$ and the CoT length filter threshold $\lambda_{c}$.
+    $\lambda_{avg}$: Average output length (tokens);
+    $\rho(\%)$: Compression rate compared to Base; $\rho_{avg}(\%)$: Average compression rate across all datasets;
+    $\gamma$ in TokenSkip: compression ratio, lower $\gamma$ means higher compression rate.
+Table 4. The CoT Compression Performance Over Different Datasets (16K Token Limit)
+Overall, SEER achieves the best performance in both reasoning efficiency and task accuracy across all three software engineering benchmarks.
+On MathQA-Python, SEER attains a pass@1 of 74.9% while reducing the average CoT length by 39.8%, indicating that substantial compression can be achieved without sacrificing correctness in code generation tasks.
+On Defect-Detection, SEER delivers the best overall performance, achieving the highest pass@1 (50.5%) together with the most aggressive compression (57.2%).
+Similarly, on Code-Search, SEER preserves the highest accuracy (77.3%) while shortening reasoning traces by 27.8%.
+These results demonstrate that SEER can effectively control CoT verbosity while maintaining, or even improving, task performance across diverse settings.
+In contrast, TokenSkip exhibits unstable behavior across compression settings.
+Under a low $\gamma$ setting (i.e., high compression rate), it performs particularly poorly: on Defect-Detection, the average CoT length even increases compared to the base model, leading to more truncation and lower accuracy.
+Although a higher $\gamma$ improves stability, TokenSkip still underperforms SEER.
+Moreover, on MathQA-Python, its pass@1 drops to around 1% regardless of the compression setting, as its token-level compression ignores code structure and produces unexecutable code.
+Other baselines that use fine-tuning methods without explicit compression operations, including Self-Trained and Naive BoN, show more stable behavior but offer limited gains.
+While these approaches achieve moderate improvements over the base model, their compression rates remain relatively low, and their accuracy consistently lags behind SEER.
+This suggests that generic fine-tuning or simple sampling strategies are insufficient to address the trade-off between reasoning length and performance.
+Prompting-based methods also achieve a certain degree of CoT length reduction, with an average compression rate of 21.9% across the three datasets.
+However, this approach comes at a noticeable cost to accuracy, as concise prompt may suppress necessary intermediate reasoning.
+On the coding task (i.e., MathQA-Python), pass@1 drops substantially by 8.1%.
+These results further reinforce our Empirical Finding 3.2.
+    Answer to RQ1: SEER consistently outperforms all baseline methods by achieving both high accuracy and substantial CoT compression across software engineering tasks.
+On average, SEER reduces CoT length by 41.6%, exceeding the strongest baseline by 10.8%, while improving pass@1 accuracy under strict context constraints.
+6.3. RQ2: Generalization Performance of SEER
+In this section, we evaluate the generalization capability of SEER across heterogeneous training domains.
+Specifically, we fine-tune the model on three different datasets and evaluate its performance on two unseen benchmarks, HumanEval and MBPP, as illustrated in Figure 4.
+Figure 4.
+Performance of SEER fine-tuned on different software engineering datasets when evaluated on HumanEval and MBPP.
+Across all software engineering fine-tuning settings, SEER consistently improves model capability by achieving higher accuracy while reducing the average length of CoT reasoning, compared to the base model.
+Figure 4 shows that fine-tuning SEER on different software engineering tasks consistently improves both task accuracy and reasoning efficiency when evaluated on unseen benchmarks.
+On HumanEval, the base model achieves about 78.0% accuracy with an average CoT length of 4,100 tokens; the fine-tuned models improve accuracy by 4.9 to 9.8 points (e.g., from 78.0% to 87.8%) while reducing the average reasoning length to around 2,400 tokens ( $\sim$40% reduction).
+On MBPP, the fine-tuned models improve accuracy by about 2.2 points (e.g., from 73.5% to 75.7%) and reduce the reasoning length to 1,680 to 2,100 tokens (up to 32% reduction).
+Overall, SEER achieves a stable CoT-length reduction on both benchmarks, indicating that it learns transferable compression behaviors.
+    Answer to RQ2: On unseen benchmarks, SEER consistently improves accuracy while reducing reasoning length.
+Accuracy gains reach up to 9.8% on HumanEval and about 2.2% on MBPP, with a reduction of roughly 30% to 40% in CoT length.
+6.4. RQ3: The Impact of Each Components
+In this section, we conduct ablation studies to isolate the contribution of each module in SEER.
+We use MathQA-Python for ablation because it represents one of the most classic and widely studied tasks in software engineering (i.e., code generation).
+We vary (1) the BoN sampling size $\mathcal{N}$ while disabling the length filter, and (2) the strictness parameter $\alpha$ in the MAD-based cutoff (which determines $\lambda_{c}$) while removing BoN.
+We additionally report the full SEER configuration that combines both components.
+All other training and evaluation settings are kept identical.
 The results are summarized in Table 5.
 Table 5. The Ablation Results of SEER on MathQA-Python Dataset
-Methods Settings Pass@1 $\mathcal{A}$ (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
+Methods Settings Pass@1 (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
 Base - 63.67 1,456 -
 Without Filter $\mathcal{N}=1$ 73.49 1,191 18.20
-$\mathcal{N}=3$ 74.45 1,177 19.16
+[…]
 $\mathcal{N}=5$ 74.40 1,163 20.12
-Without BoN $\lambda_{c}=2\times\lambda_{ada}$ 74.40 1,071 26.44
-$\lambda_{c}=1.5\times\lambda_{ada}$ 74.19 918 36.95
-$\lambda_{c}=\lambda_{ada}$ 73.50 878 39.70
-$\lambda_{c}=0.5\times\lambda_{ada}$ 72.76 836 42.58
+Without BoN $\alpha=2$ 74.40 1,071 26.44
+[…]
+$\alpha=1$ 73.50 878 39.70
+[…]
 SEER
-$\mathcal{N}=3$ $\lambda_{c}=\lambda_{ada}$
+$\mathcal{N}=3$ $\alpha=1$
 74.88 877 39.77
-When only BoN sampling is applied without the filter, increasing $\mathcal{N}$ from 1 to 3 improves accuracy from 73.49% to 74.45%, while the average CoT length decreases slightly from 1,191 to 1,177 tokens.
-However, further increasing $\mathcal{N}$ to 5 does not provide additional benefit and even slightly reduces accuracy (74.40%), indicating that the advantage of BoN sampling quickly saturates.
-In this setting, the compression rate remains relatively modest, ranging from 18.20% to 20.12%.
-In contrast, applying only the adaptive filter leads to a substantial increase in compression.
-For instance, when $\lambda_{c}=2\times\lambda_{ada}$, the compression rate reaches 26.44%, and as the threshold becomes stricter ($\lambda_{c}=0.5\times\lambda_{ada}$), compression improves further to 42.58%.
-Nevertheless, overly aggressive filtering reduces accuracy, dropping from 74.40% to 72.76%.
-These results show that while the filter is more effective than BoN in reducing the redundancy in CoT, it risks discarding essential reasoning steps when applied too strongly.
-To balance accuracy and compression rate, we set $\mathcal{N}=3$ and $\lambda_{c}=\lambda_{ada}$ in our method.
-When both components are combined in SEER with these settings, the method achieves the best trade-off, with the highest accuracy of 74.88% and a competitive compression rate of 39.77%.
-This demonstrates that the two modules are complementary: BoN sampling enhances robustness, while the adaptive filter enforces compactness, and together they enable SEER to deliver concise yet accurate reasoning.
-[…]
-Answer to RQ2: The ablation results reveal that BoN sampling alone offers only limited gains, while the adaptive filter strongly enhances compression but may harm accuracy if too strict.
-Their combination achieves the best trade-off, yielding both higher accuracy and effective compression, highlighting the complementary roles of the two components.
-6.4. RQ3: Generalizability of SEER
-In this section, we evaluate the generalizability of SEER by fine-tuning it on one dataset and testing on the remaining three datasets. The results are presented in Table 4.
-Table 4. The Generalizability Evaluation Result of SEER Across All Tasks
-MathQA-Python Defect-Detection Code-Search GSM-8K
-Models $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$
-Base 63.7 1456 - 44.7 1,836 - 72.4 472 - 90.1 1,309 -
-MathQA-Py Fine-tune 74.9 877 39.8 47.1 1131 38.4 77.3 395 16.3 87.2 748 42.9
-Defect-Detection Fine-tune 64.2 1279 12.2 50.5 785 57.2 72.2 339 28.2 90.4 1128 13.8
-Code-Search Fine-tune 64.1 1104 24.2 51.4 672 63.4 77.3 362 23.3 89.8 853 34.8
-GSM-8K Fine-tune 62.5 1138 21.8 48.5 1483 19.2 79.6 435 7.8 90.5 677 48.3
-Overall, the findings demonstrate that SEER exhibits strong cross-task generalization.
-Regardless of which dataset is used for fine-tuning, the resulting models generally achieve higher accuracy compared with the base model.
-Only a few cases show a slight decrease, such as fine-tuning on GSM-8K and then evaluating on MathQA-Python (62.5% vs.
-63.7%).
-In all cases, the compression rate remains consistently above zero.
-For example, fine-tuning on MathQA-Python compresses average reasoning length on GSM-8K from 1,309 tokens to 748 tokens (42.9% compression), while still preserving high accuracy (87.2%).
-This indicates that SEER is able to reduce reasoning length while maintaining competitive accuracy even on out-of-domain tasks.
-Interestingly, in some settings, the fine-tuned models even surpass those trained directly on their target dataset.
-For instance, the model fine-tuned on Code-Search achieves 51.4% accuracy on Defect-Detection, which is higher than the model fine-tuned on Defect-Detection (50.5%).
-These results suggest that SEER captures reusable reasoning patterns that extend beyond the dataset it was trained on, rather than overfitting to a single task.
-Figure 4.
-Performance of SEER fine-tuned on different datasets when evaluated on HumanEval.
-Fine-tuning on coding-related datasets such as MathQA-Python and Code-Search yields the best generalization with shorter CoT lengths, while non-coding datasets like GSM-8K still provide moderate improvements over the base model.
-To further assess its generalizability, we evaluate the four fine-tuned models on the HumanEval dataset, as shown in Figure 4.
-The results reveal two key insights.
-First, all fine-tuned models achieve noticeable improvements in both accuracy and CoT length over the base model, confirming that the knowledge learned from different tasks generalizes to program synthesis.
-Second, the reduction in reasoning length remains consistent, suggesting that SEER’s compression mechanism is stable even when applied to tasks outside the training domain.
-This demonstrates that the framework does not merely adapt to a specific dataset but instead learns reasoning strategies that generalize across diverse software engineering and mathematical tasks.
-[…]
-Answer to RQ3: SEER demonstrates strong cross-task generalization.
-Even when evaluated on out-of-domain datasets, it consistently improves accuracy while reducing CoT length.
-Fine-tuning on coding datasets further generalizes well to HumanEval, confirming that SEER learns reusable reasoning strategies beyond its training domain.
-6.5. RQ4: The Ability of Reducing Reasoning Loops
+In the Without Filter setting (Table 5), we vary the BoN sampling size $\mathcal{N}$.
+Increasing $\mathcal{N}$ from 1 to 3 improves accuracy from 73.49% to 74.45%, while the average CoT length decreases slightly from 1,191 to 1,177 tokens.
+However, further increasing $\mathcal{N}$ to 5 does not provide additional benefit and slightly reduces accuracy (74.40%), indicating that the effect of BoN quickly saturates.
+In this setting, the compression rate remains modest (18.20% to 20.12%).
+In the Without BoN setting (Table 5), we vary the strictness parameter $\alpha$ in our MAD-based cutoff $\lambda_{\mathrm{mad}}(\alpha)=\tilde{\lambda}+\alpha\cdot\mathrm{MAD}$.
+As the filter becomes stricter (from $\alpha=2$ to $\alpha=0.5$), the compression rate increases from 26.44% to 42.58%, but accuracy drops from 74.40% to 72.76%.
+This shows a clear trade-off: stronger filtering yields higher compression but can remove useful reasoning information.
+We use $\mathcal{N}=3$ and $\alpha=1$ as the default setting in SEER.
+With both components enabled, SEER achieves the best balance, reaching the highest accuracy (74.88%) with a competitive compression rate (39.77%).
+    Answer to RQ3: BoN alone provides only modest gains (e.g., accuracy increases from 73.49% to 74.45% when $\mathcal{N}$ increases from 1 to 3, with compression of 18.20% to 20.12%).
+Making the length filter stricter increases compression (from 26.44% to 42.58%) but reduces accuracy (from 74.40% to 72.76%).
+Using both components together achieves the best balance, reaching 74.88% accuracy with 39.77% compression.
+6.5. RQ4: Effectiveness in Mitigating Reasoning Loops
 Infinite reasoning loops occur when a model repeatedly generates identical reasoning segments until the output reaches the maximum context length and is truncated.
 This phenomenon wastes computational resources, inflates latency, and often produces incomplete or invalid final answers.
-Because such loops can arise unpredictably across models of different sizes, they represent a critical robustness concern.
+Since such loops can arise unpredictably across models of different sizes, they represent a critical robustness concern.
 In this section, we investigate whether our approach can mitigate this problem.
-For each dataset, we adapt our method to fine-tune a model specific to that dataset and evaluate it against the base model.
-The results in Table 5 show that SEER substantially reduces reasoning loops in all cases.
-On MathQA-Python, loop occurrences drop from 85 to 23 (72.9% reduction).
-On Code-Search, the count falls from 15 to 1 (93.3% reduction), effectively eliminating loops.
-The largest improvement appears in Defect-Detection, where loops decrease from 222 to just 7 (96.8% reduction).
-Even on GSM8K, which generally contains shorter reasoning chains, loops drop from 54 to 6 (88.9% reduction).
-Table 5. The Performance of SEER in Reducing Truncation and Reasoning Loops
-MathQA-Python Code-Search Defect-Detection GSM8K
-Truncation Loops Truncation Loops Truncation Loops Truncation Loops
-Base 110 85 18 15 235 222 59 54
-SEER 30 23 1 1 7 7 8 6
-During our empirical analysis in Section 3, we find that most truncations in the base model are directly caused by such loops.
+We evaluate SEER on all three datasets, consistent with RQ1.
+For each dataset, we fine-tune a dataset-specific model using our method and compare it against the base model.
+Table 6 summarizes the results.
+SEER substantially reduces reasoning loops in all cases, e.g., from 85 to 23 on MathQA-Python (72.9%), from 15 to 1 on Code-Search (93.3%), and from 222 to 7 on Defect-Detection (96.8%).
+Table 6. The Effectiveness of SEER in Mitigating Truncation and Reasoning Loops
+MathQA-Python (NN=1883) Code-Search (NN=9604) Defect-Detection (NN=2732)
+Truncation # (%) Loops # (%) Truncation # (%) Loops # (%) Truncation # (%) Loops # (%)
+Base 110 (5.8%) 85 (4.5%) 18 (0.19%) 15 (0.16%) 235 (8.6%) 222 (8.1%)
+[…]
+In our Empirical Findings 3.2, we find that most truncations in the base model are directly caused by such loops.
 This behavior wastes much of the available context budget and prevents the model from completing its reasoning.
-Our adaptive filtering strategy compresses CoT length, which mitigates loop formation and reduces truncation, as shorter reasoning provides fewer chances for repetitive output.
-This dual effect, loop suppression and truncation reduction, further enhances the robustness and practicality of our method in real-world deployments.
+Our adaptive filtering strategy reduces CoT length, mitigates loop formation and reduces truncation, as shorter reasoning provides fewer chances for repetitive output.
 Overall, these results confirm that SEER offers a robust defense against infinite reasoning loops, complementing its primary goal of efficient CoT compression and providing benefits for latency, reliability, and output quality.
-[…]
     Answer to RQ4: SEER effectively suppresses infinite reasoning loops across all datasets, with reductions ranging from 72.9% on MathQA-Python to 96.8% on Defect-Detection.
-Because most truncations in the base model stem from such loops, the adaptive filtering strategy not only prevents pathological repetition but also lowers truncation frequency.
+Because most truncations in the base model stem from such loops, the adaptive filtering strategy not only prevents looping problems but also lowers truncation frequency.
 7. Discussion
@@
 An important aspect of our framework is its adaptability to different fine-tuning paradigms.
-Beyond full-parameter supervised fine-tuning (SFT), we also explored parameter-efficient alternatives such as LoRA.
+Beyond full-parameter supervised fine-tuning (SFT), we also explore parameter-efficient alternatives such as LoRA.
 The comparison reveals two insights.
@@
 Table 7. The Performance Comparison Between Using SFT and PEFT Fine-tuning Method
-Pass@1 $\mathcal{A}$ (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
+$pass@1(\%)$ Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
 Base 63.67 1,456 -
-PEFT (LoRA) 67.71 949 34.8
+[…]
 SFT 74.88 877 39.8
@@
 7.2.1. Internal Validity
-Internal validity concerns whether the observed improvements of SEER truly arise from our proposed framework rather than confounding factors.
-One potential threat is the reliance on a single model family, namely the DeepSeek-R1-Distill-Qwen-7B.
-Although we carefully fine-tuned and evaluated under consistent settings to reduce bias, the conclusions may partially reflect properties of this architecture rather than general effectiveness.
-To mitigate this risk, we performed ablation studies (varying BoN sampling and adaptive filtering thresholds) to confirm that the observed gains stem from the specific design choices of SEER rather than incidental artifacts.
-Future work will extend the evaluation to other model families, including implicit CoT models that do not explicitly output reasoning traces.
-Moreover, testing on closed-source models with advanced reasoning capabilities (e.g., OpenAI O1, Gemini, Claude’s Thinking series) will help validate whether the improvements are architecture-independent.
+Internal validity concerns whether the observed improvements of SEER truly arise from the proposed framework rather than confounding factors.
+One potential threat is the reliance on a single model family, namely DeepSeek-R1-Distill-Qwen-7B.
+Although all methods are fine-tuned and evaluated under consistent settings, some effects may reflect properties of this architecture.
+To mitigate this risk, we conduct ablation studies by varying BoN sampling and adaptive filtering thresholds, confirming that the gains are attributable to the core design of SEER rather than incidental artifacts.
+Another potential concern is filtering bias.
+Our filter may exclude some difficult instances that naturally require longer CoTs.
+This is partially intended: SEER does not aim to create reasoning ability from scratch, but to encourage an already reasoning-capable model to produce concise, non-redundant reasoning.
+Since extremely long CoTs are often dominated by overthinking or looping, filtering these traces helps the model learn efficient reasoning patterns.
+To avoid removing problems entirely, filtering is applied to CoT traces rather than problem instances via BoN sampling.
+More importantly, our method does not incur a pass@1 loss after filtering, suggesting that this filtering strategy helps the model learn to generate reasoning that is both concise and effective.
 7.2.2. External Validity
-External validity concerns the generalizability of our findings beyond the datasets and settings considered in this work.
-Our experiments primarily focus on software engineering tasks, as they represent one of the most critical application domains for reasoning models.
-To mitigate the risk of over-specialization, we additionally included a mathematical reasoning dataset (GSM-8K), which provides a different reasoning style and serves as an initial step toward broader validation.
-Nevertheless, important domains such as commonsense reasoning, legal QA, and scientific problem solving remain unexplored.
-Moreover, all evaluations were conducted with fixed prompts and prompting strategies.
-While our method is designed to be prompt-agnostic, alternative instruction formats, few-shot configurations, or domain-specific prompts could influence the observed trade-off between compression and accuracy.
-Finally, although the datasets used contain tasks of varying difficulty, we have not systematically examined whether compression benefits differ across problem complexity levels.
-To further strengthen external validity, we plan to expand our evaluation to additional reasoning domains and more diverse datasets, as well as investigate robustness under alternative prompting strategies.
-Future work will also include stratified analyses by task difficulty to better understand whether SEER is equally effective on both simple and challenging instances.
+External validity concerns whether our findings generalize beyond the evaluated settings.
+A potential limitation is that our study focuses on software engineering tasks, which may restrict applicability to other reasoning domains.
+We mitigate this threat by evaluating SEER on three representative software engineering benchmarks covering distinct task types, including code generation, defect detection, and code search, each involving different reasoning patterns.
+Another concern is sensitivity to prompt design.
+To reduce this risk, all methods are evaluated using fixed prompts and consistent prompting strategies, ensuring that observed improvements are attributable to the proposed method rather than prompt-specific effects.
+Moreover, SEER is designed to be largely prompt-agnostic, further limiting reliance on particular instruction formats.
+Finally, the effectiveness of CoT compression may vary with problem difficulty.
+Although we do not explicitly stratify instances by complexity, the evaluated datasets naturally contain problems of varying difficulty, and results are reported over full benchmarks rather than curated subsets.
+A finer-grained analysis by difficulty is left for future work.
 8. Related Works
@@
 Token-Budget (Han et al., 2024) proposes a token-budget-aware reasoning framework that dynamically adjusts reasoning length based on estimated complexity.
-CoD (Xu et al., 2025a) limits reasoning to within five words via a “Chain-of-Draft” template.
+CoD (Xu et al., 2025b) limits reasoning to within five words via a “Chain-of-Draft” template.
 Token Complexity (Lee et al., 2025) systematically analyzes prompt compression strategies, highlighting that many still operate far from theoretical efficiency limits.
@@
 SCoT (Li et al., 2025) was among the first to apply structured CoT to code generation.
-The emergence of Deepseek-R1 (DeepSeek-AI, 2025b) further highlighted the value of streamlined CoT for software engineering tasks.
+The emergence of DeepSeek-R1 (DeepSeek-AI, 2025) further highlighted the value of streamlined CoT for software engineering tasks.
 Liu et al.
@@
 CoD (Yang, 2025) adapts the “Chain-of-Draft” approach to software engineering tasks, greatly reducing token usage.
-However, existing SE-focused methods often target specific task scenarios or rely on inference-time control, which may not eliminate pathological failure modes, such as infinite reasoning loops.
+However, existing SE-focused methods often target specific task scenarios or rely on inference-time control, which may fail to eliminate severe failure modes such as looping behaviors.
 Our method differs by explicitly filtering out redundant or looping reasoning during training, allowing the model to internalize concise, stable reasoning behaviors.
-This makes it robust not only in software engineering tasks but also in mathematical reasoning, while retaining interpretability and avoiding the need for task-specific prompt engineering.
+This makes it robust in software engineering tasks while retaining interpretability and avoiding the need for task-specific prompt engineering.
 9. Conclusion
-Our work systematically examined the inefficiency of excessively long CoT reasoning, identifying that long CoT can lead to truncation and overthinking, resulting in decreased accuracy.
-We then present SEER, a self-enhancing framework for adaptive CoT compression.
-Through extensive experiments on software engineering and mathematical reasoning tasks, SEER reduces CoT length by an average of 42.1% while maintaining or even improving accuracy.
-Our ablation results confirm the complementary roles of BoN sampling and adaptive filtering, and cross-domain evaluations demonstrate SEER’s strong generalizability.
-SEER also effectively mitigates truncation and suppresses infinite reasoning loops, addressing key efficiency and robustness issues.
-These findings highlight SEER as a practical approach for making CoT-enhanced LLMs more efficient and reliable.
-Future work will extend SEER to broader reasoning domains, apply it to diverse model architectures, and explore integration into latency-sensitive real-world systems.
+In this paper, we investigate the practical costs of overly long CoT reasoning in software engineering settings.
+We find that many modern reasoning models produce excessively verbose CoTs, which substantially increases token cost and can trigger truncation; moreover, many truncations are caused by looping behaviors.
+To address these challenges, we proposed SEER, a self-enhancing framework for adaptive CoT compression that learns concise reasoning patterns from self-generated outputs by combining BoN sampling with a lightweight, data-driven CoT length filter.
+Across three software engineering tasks, SEER reduces CoT length by an average of 41.6% while improving performance.
+Cross-domain evaluations indicate good generalization to unseen benchmarks.
+Overall, SEER makes CoT-enhanced LLMs more efficient and robust under real-world context constraints.
+Future work includes extending SEER to more domains and model families and integrating it into latency-sensitive agent pipelines.